### v2

# Conformal Tradeoffs: Guarantees Beyond Coverage

Abstract
Deployed conformal predictors are long-lived decision infrastructure reused over finite operational windows. In practice, stakeholders care not only about marginal coverage, but also about operational quantities: how often the system commits versus defers, and what error exposure it induces when it acts. These deployment-facing quantities are not determined by coverage alone: identical calibrated thresholds can yield markedly different operational profiles depending on score geometry. We develop tools for operational certification and planning beyond coverage for split conformal prediction. First, Small-Sample Beta Correction (SSBC) inverts the exact finite-sample rank/Beta law to map a user request $(\alpha^{\star},\delta)$ to a concrete calibration grid point with PAC-style semantics, yielding explicit finite-window coverage guarantees for a reused deployed rule. Second, because no distribution-free pivot exists beyond coverage, we propose Calibrate-and-Audit: an independent audit set supports certified finite-window predictive envelopes (Binomial/Beta-Binomial) for key operational quantities—commitment frequency, deferral, and decisive error exposure—and related metrics via linear projection, without committing to a scalar objective. Third, we give a geometric characterization of the feasibility constraints and regime boundaries induced by a fixed conformal partition, clarifying why operational quantities are coupled and how calibration navigation trades them off. The result is an operational menu rthat traces attainable operational profiles (Pareto trade-offs) and attach finite-window uncertainty envelopes to each regime. We illustrate the approach on benchmark molecular toxicity and aqueous solubility datasets.
1 Introduction: deployment-facing conformal prediction
Many classifiers are deployed as long-lived decision infrastructure rather than one-off prediction engines. A single trained scoring model is integrated into downstream workflows and reused over finite operational windows. In this regime, the deployed object is not the score function alone, but the combined rule tying together the scoring model, the one-time calibration procedure, and the downstream action convention. The observables that matter to stakeholders are finite-window operational rates: how often the system commits versus defers/abstains, and conditional on commitment the decisive error exposure and purity induced by attached actions (cf. rejection options and selective classification (Chow, 1970)). These realized rates drive resource planning, safety and compliance, and near-term impact.
Conformal prediction (Vovk et al., 2005; Shafer and Vovk, 2008) is a natural starting point for deployment-facing calibration: it sets score thresholds on held-out data and yields finite-sample, distribution-free coverage under exchangeability (Papadopoulos et al., 2002). Coverage alone, however, is not an operational specification that captures the full range of deployment-facing quantities (Hullman et al., 2025). Once a deployment convention is fixed (e.g., “commit on singletons, defer on hedges or empty sets”), the realized operational profile depends on how calibration choices partition score space and how mass and labels populate the induced regions. As a result, two conformal rules with identical marginal coverage can behave very differently in deployment.
1.1 The gap: coverage does not determine operational behavior
Standard conformal guarantees that the true label lies in the prediction set at the target coverage rate (under exchangeability). But coverage alone does not determine the operational profile of a deployed conformal rule. In particular, different calibrated predictors with the same nominal coverage can exhibit markedly different operational characteristics. Examples include:
  • Commitment vs. deferral: the frequency of singleton outputs (and thus decisive actions) versus hedges/abstentions.
  • Decisive error exposure: the error rate among committed predictions, which is the error that “escapes” deferral mechanisms.
  • Trade-off coupling: improving one operational rate (e.g., reducing deferral) may force degradation in another (e.g., increasing decisive error), with couplings determined by the score distribution geometry rather than by coverage alone.
These are first-order questions for operational readiness and trust calibration, yet they are not easily answered by coverage certificates alone.
1.2 Our approach: calibrate, expose geometry, and audit once
We study a single deployed split-conformal rule through a calibration-conditional, certification-style lens. A calibration choice fixes cutpoints on the score axis, inducing a finite partition of score/output space into regions with set-valued outcomes (e.g., singleton, hedge, abstain). The key device in this paper is to make this intermediate geometry explicit and auditable: once the region partition is fixed, many policy- and Key Performance Indicator (KPI)-specific rates are obtained by simple projection from the induced joint region–class label distribution. Our development focuses on binary classification. Even in this simplest nontrivial setting, the induced geometry already reveals the main operational trade-offs. Within this setting, we can study trade-offs across calibration settings and derive distribution-free uncertainty statements about future realized finite-window rates for the deployed rule.
Figure 1 previews the main object produced by this paper: the map from calibration choices to the induced joint region–class label distribution. This representation can then be interrogated through Pareto analysis to expose operational trade-offs before a price point or consequence cost has been specified.
Figure 1: Operational menu induced by a calibrated conformal rule on a synthetic model. Left: a calibration choice fixes score cutpoints that partition score space into set-valued outcomes (e.g., singleton, hedge, abstain) and associated region–class label probabilities. A deployment convention determines which outcomes trigger commitment versus deferral. Right: sweeping calibration settings traces the attainable set of selected operational rates; the oriented Pareto frontier highlights nondominated operating regimes.
1.3 Contributions
This paper studies operational quantities as explicit objects of exploration and certification for deployed conformal predictors. Our contributions are:
(1) Coverage semantics via Small-Sample Beta Correction (SSBC).
We invert the exact finite-sample rank/Beta law for split conformal (Marques, 2025; Angelopoulos and Bates, 2023) to map a user request $(\alpha^{\star},\delta)$ (e.g., “at least 90% coverage with at least 90% confidence”) to the least conservative split-conformal grid point satisfying a Probabily Approximately Correct (PAC) tail constraint:
$\mathbb{P}_{\mathcal{D}_{\mathrm{cal}}}\!\bigg(\mathbb{P}\!\big(Y\in\hat{C}(X)\mid\mathcal{D}_{\mathrm{cal}}\big)\geq 1-\alpha^{\star}\bigg)\geq 1-\delta.$
This provides an explicit, auditable finite-sample coverage guarantee for the deployed rule and serves as a semantic anchor for subsequent operational navigation. In the binary class-conditional case, SSBC collapses a four-dimensional user specification into a two-dimensional calibration coordinate, simplifying trade-off exploration.
(2) Calibrate-and-Audit for operational certification beyond coverage.
Operational rates such as commitment, deferral, and decisive error exposure are not rank-determined like coverage and therefore do not admit a conformal “coverage-style” pivot that determines their distribution. We therefore introduce a two-stage design for studying these quantities:
  • Calibrate: fix thresholds on $\mathcal{D}_{\mathrm{cal}}$, inducing a region partition $R_{\tau}$.
  • Audit: use an independent exchangeable set $\mathcal{D}_{\mathrm{audit}}$ to estimate the joint region–class label table $\{p_{r,y}\}$.
The joint table is a reusable summary object: many operational KPIs can be written as linear projections (or monotone transforms) of $\{p_{r,y}\}$. Together with the induced Binomial/Beta–Binomial sampling model for events of interest, it provides finite-window predictive envelopes for future realized operational rates and a way to compare calibration settings without re-auditing. We also provide a conservative leave-one-out proxy for settings where there is no budget for an independent audit set (Appendix D).
(3) Geometry and feasibility of operational trade-offs.
We expose how a fixed conformal partition couples attainable operational rates. In the binary probability-normalized case:
  • Regime boundaries: the sum $\tau_{0}+\tau_{1}$ determines whether the system can hedge ($\tau_{0}+\tau_{1}>1$) or must reject ($\tau_{0}+\tau_{1}<1$), inducing sharp transitions in attainable profiles (Section 4).
  • Conservation constraints: varying thresholds reallocates mass across regions, explaining structural trade-offs between commitment and decisive error exposure.
  • Cost coherence: we derive compatibility conditions between region-based action conventions and scalar cost models, in advance of any cost model specification that are linear transformations of the auditable primitives (Appendix F).
Together, these results clarify why some operational profiles are infeasible and help characterize the trade-offs induced by conformal calibration choices.
1.4 Positioning and related work
Conformal prediction constructs set-valued predictors with finite-sample coverage guarantees under exchangeability (Vovk et al., 2005; Shafer and Vovk, 2008; Angelopoulos and Bates, 2023). Conformal risk control (CRC) generalizes this perspective by choosing thresholds to satisfy user-specified scalar risk constraints under related assumptions (Angelopoulos et al., 2024; Bates et al., 2021), and extensions address covariate shift via reweighting or adaptation (Tibshirani et al., 2019; Fannjiang et al., 2022). Related work also studies stronger conditional notions of validity beyond marginal coverage, including training-conditional and other conditional guarantees (Bian and Barber, 2023; Gibbs et al., 2025). A growing literature connects conformal methods to downstream decision-making: conformal decision theory calibrates decision rules (often online) to control realized risk (Lekeufack et al., 2023), while decision-theoretic foundations motivate conformal sets as uncertainty objects for risk-averse agents (Kiyani et al., 2025).
Recent Human-Computer Interaction (HCI)-oriented work emphasizes a complementary concern: conformal sets can be a “murky” interface for human decision-making, because validity and common efficiency summaries do not identify the action-relevant consequences users care about (Hullman et al., 2025). Fig. 1 of our paper makes one such concern concrete: each conformal set corresponds to a region in the calibration-induced partition and is associated with specific region-conditional class frequencies, which are not apparent from set membership alone.
Closest in spirit to our operational menu viewpoint is the inverse CRC of Zhou and Zhu (2025), who trace certified miscoverage–regret trade-offs across robustness levels for predict-then-optimize pipelines, auditing coverage and regret as the downstream objects of interest. A different line of work targets efficiency by selecting, from a family of candidate models, the conformal predictor that minimizes a set-size functional (e.g., interval length/volume, or prediction-set cardinality in classification) subject to validity constraints (Yang and Kuchibhotla, 2021). This is well-motivated when downstream utility is monotone in set size, but our objective differs: in many deployments, smaller cardinality does not necessarily imply higher utility, nor lower uncertainty in induced region-conditional label frequencies. In particular, aggressive singleton behavior can reduce average set size while increasing decisive error exposure and degrading the operational profile that stakeholders ultimately care about. When organizational priorities are still being determined, set size can therefore be an uninformative proxy for decision quality.
Our focus is complementary to these lines of work. Rather than guaranteeing coverage or a single scalar risk functional, we certify finite-window operational rate vectors for a fixed deployed rule and support multi-objective exploration without committing to a scalar objective. The audited primitive is the intermediate geometry induced by calibration—namely the region partition and its joint region–class label table—from which many deployment policies and KPIs follow by projection. This calibrate-and-audit scheme also clarifies the internal statistical structure of the prediction sets by making explicit how conformal calibration induces the partition from which downstream quantities are derived.
1.5 Roadmap
Section 2 formalizes the calibration-conditional viewpoint for a fixed deployed object and defines the region–label table as the auditable primitive. Section 3 presents SSBC and the Calibrate-and-Audit framework, yielding certified finite-window envelopes for operational rates and Pareto-based trade-off navigation. Section 4 characterizes regime boundaries and feasibility constraints induced by the conformal partition. Section 5 validates the framework empirically on Tox21 toxicity prediction and aqueous solubility prediction, and Section 6 discusses limitations and extensions.
2 Setting and notation: calibration-conditional viewpoint
We study a single deployed classifier produced by (i) training a scoring model on a training set and then treating it as fixed, followed by (ii) a one-time split conformal calibration on an exchangeable calibration sample. Accordingly, all randomness in our analysis is over the one-time calibration draw and future deployment data; we do not analyze variability due to retraining or repeated recalibration. Later, to make operational rates beyond coverage auditable with certified finite-sample guarantees, we introduce a separate exchangeable audit set held out alongside the calibration sample.
2.1 Data splits and exchangeability assumptions
Let $\mathcal{D}_{\mathrm{train}}$ denote a training dataset used to fit a base probabilistic classifier (Goodfellow et al., 2016). After training, the resulting score function is treated as fixed.
Let
$\mathcal{D}_{\mathrm{cal}}=\{(X_{i}^{c},Y_{i}^{c})\}_{i=1}^{n_{\mathrm{cal}}},\qquad\mathcal{D}_{\mathrm{audit}}=\{(X_{i}^{a},Y_{i}^{a})\}_{i=1}^{n_{\mathrm{audit}}}.$
The calibration sample $\mathcal{D}_{\mathrm{cal}}$ is used once to set the deployed rule’s thresholds; the audit sample $\mathcal{D}_{\mathrm{audit}}$ is held out and reserved for measuring operational quantities after thresholds are fixed, and is not used to choose thresholds. This use of a double calibration set is not without precedent; similar split designs appear in Yang and Kuchibhotla (2021).
We assume calibration, audit, and future deployment examples are jointly exchangeable (conditional on the fixed trained scoring model). This is the standard split conformal assumption that yields distribution-free finite-sample validity for coverage, and it is the only distributional condition used throughout. All guarantees are scoped to this exchangeable setting; we do not consider covariate or label shift in this paper. In this work we focus on binary classification to isolate the core principles and trade-offs in the clearest setting; extensions to richer output spaces are left for future work.
2.2 Scores and calibration thresholds
Let $\mathcal{Y}=\{1,\ldots,K\}$, and let $s:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ denote a nonconformity (score) function (Shafer and Vovk, 2008; Lei et al., 2018). A common choice for probabilistic-like classifiers is $s(x,y)=1-P(y\mid x)$. While this specific form is not required, it is used in the examples throughout the paper.
Given $\mathcal{D}_{\mathrm{cal}}$, compute calibration nonconformity scores
$S_{i}:=s(X_{i}^{c},Y_{i}^{c}),\qquad i=1,\ldots,n_{\mathrm{cal}},$
and let $S_{(1)}\leq\cdots\leq S_{(n_{\mathrm{cal}})}$ denote the sorted values. Split conformal calibration selects a threshold as an order statistic (Papadopoulos et al., 2002; Vovk, 2012b)
$\tau:=S_{(k)},$
where the index $k\in\{1,\ldots,n_{\mathrm{cal}}\}$ is determined by the desired miscoverage level $\alpha^{\star}$ (up to deterministic tie-breaking). It is often convenient to parameterize the same grid by the miscoverage index $u:=n_{\mathrm{cal}}+1-k$, corresponding to the discrete miscoverage level $\alpha_{\mathrm{grid}}=u/(n_{\mathrm{cal}}+1)$.
We use class-conditional split conformal throughout (Vovk, 2012b; a), where thresholds $\tau_{y}$ are computed separately from calibration scores restricted to $Y_{i}=y$. The pooled split conformal convention is recovered as the special case with shared thresholds across classes, i.e., $\tau_{y}\equiv\tau$ for all $y$.
Thus, for a given $(\mathcal{D}_{\mathrm{cal}},\alpha^{\star})$, calibration produces a vector of class-conditional thresholds $\tau=(\tau_{1},\ldots,\tau_{K})$ that is held fixed during deployment.
2.3 Regions, observability, and policies
Once thresholds are fixed, they induce a finite discrete region by recording which threshold inequalities are satisfied. In the multiclass case it is convenient to write the score vector
$s(x):=(s(x,1),\ldots,s(x,K))$
and class-conditional thresholds
$\tau:=(\tau_{1},\ldots,\tau_{K}),$
with $\tau_{y}\equiv\tau$ in the pooled case. The region map is defined by
$R_{\tau}(x):=\big(\mathbf{1}\{s(x,1)\leq\tau_{1}\},\ldots,\mathbf{1}\{s(x,K)\leq\tau_{K}\}\big)\in\{0,1\}^{K}.$
Thus calibration fixes a region map $R_{\tau}:\mathcal{X}\to\{0,1\}^{K}$, a finite partition of score space used throughout deployment. The map $R_{\tau}$ is determined entirely by $\tau$. However, the induced distribution of class labels $Y$ given the region $R_{\tau}(X)$, i.e. which regions carry mass and how much, depends on the deployment data-generating process. This discrete region partitioning is our decision and exploration interface.
Figure 2: Region definitions and region mass under different data regimes. (A) Thresholds $\tau_{0},\tau_{1}$ induce a finite region partition $R_{\tau}(x)$ independent of data and policy. (B–C) Under probability-normalized scores, data support is restricted to the diagonal, determining which regions carry nonzero mass for different threshold configurations. (D) With unconstrained scores, mass is distributed across a score-manifold, not solely no the diagonal. No deployment policy is applied in this figure.
Operationally, deployment is observed through the joint process $(R_{\tau}(X),Y)$. All quantities considered in this paper are functions of region frequencies and region-conditioned label composition under fixed thresholds. We treat region identity as the auditable primitive; within-region heterogeneity is intentionally not modeled in this framework. Accordingly, observations that share the same realized region label $R_{\tau}(X)=r$ are treated as exchangeable for the purpose of estimation and certification.
To formally link regions to downstream consequences, we introduce the concept of a deployment policy. A deployment policy $\pi$ maps a realized region $R_{\tau}(x)$ to a reported output, such as a (possibly empty) set of labels—with abstention represented by $\varnothing$—or a single committed label. The deployed predictor is written as
$\hat{C}_{\pi}(x):=\pi(R_{\tau}(x)).$
The standard split conformal predictor corresponds to the set inclusion policy:
$\pi_{\mathrm{SI}}(r):=\{y\in\mathcal{Y}:r_{y}=1\},$
yielding
$\hat{C}_{\pi_{\mathrm{SI}}}(x)=\{y\in\mathcal{Y}:s(x,y)\leq\tau_{y}\},$
with $\tau_{y}\equiv\tau$ in the pooled case.
The motivation for introducing policies is to separate calibration from action. Calibration fixes thresholds and therefore a finite set of region outcomes; deployment then chooses how to act on those outcomes. Writing $\hat{C}_{\pi}(x)=\pi(R_{\tau}(x))$ lets us swap or optimize decision rules for operational objectives (commit/defer/reject) without rewriting the calibration step, and it provides a common language for auditing and comparing policies.
As an example beyond set inclusion, consider a region-triggered commit policy. In the binary case $\mathcal{Y}=\{0,1\}$, let $A\subseteq\{0,1\}^{2}$ and $a\in\{0,1\}$, and define
$\pi_{A\to a}(r)=\begin{cases}\{a\},&r\in A,\\ \varnothing,&r\notin A.\end{cases}$
In words, the policy commits to action $a$ only when the realized conformal region belongs to the pre-specified trigger set $A$; otherwise it abstains. Thus $A$ acts as an explicit commitment gate on the fixed calibrated region partition.
Figure 3 visualizes several other deployment policies acting as deterministic projections on a fixed region structure.
Figure 3: Deployment policies as projections on a fixed region structure. Panel A shows the region partition $R_{\tau}(x)$. Panels B–D apply different region-based deployment policies $\pi$ to the same regions and data support, yielding reported outputs $\hat{C}_{\pi}(x)=\pi(R_{\tau}(x))$. Differences between panels arise solely from the choice of policy.
2.4 Auditing primitives: region–label tables and policy projections
Conditional on $\mathcal{D}_{\mathrm{cal}}$, the deployed rule is fixed, and the primitive observable for auditing is the pair $(R_{\tau(\theta)}(X),Y)$ on an exchangeable labeled split. Here $\theta\in\Theta$ denotes a calibration setting that determines thresholds $\tau(\theta)$ and hence the region map $R_{\tau(\theta)}$. Two representations are useful: the full region–class label table, and operational indicators obtained as its projections.
(A) Region–class label joint table.
Define the calibration-conditional joint probabilities
$p_{r,y}(\theta):=\mathbb{P}\!\big(R_{\tau(\theta)}(X)=r,\ Y=y\mid\mathcal{D}_{\mathrm{cal}}\big),\qquad(r,y)\in\mathcal{R}\times\mathcal{Y}.$
These are fixed (but unknown) constants for the deployed rule. The audit set is used to estimate them via the counts
$K^{\mathrm{audit}}_{r,y}(\theta):=\sum_{i=1}^{n_{\mathrm{audit}}}\mathbf{1}\{R_{\tau(\theta)}(X_{i}^{a})=r,\ Y_{i}^{a}=y\},\qquad\widehat{p}^{\,\mathrm{audit}}_{r,y}(\theta):=\frac{K^{\mathrm{audit}}_{r,y}(\theta)}{n_{\mathrm{audit}}}.$
The region map $R_{\tau(\theta)}$ is fixed from $\mathcal{D}_{\mathrm{cal}}$, while the counts $K^{\mathrm{audit}}_{r,y}(\theta)$ are computed on $\mathcal{D}_{\mathrm{audit}}$ only.
Because $R_{\tau(\theta)}$ is a finite partition of score space, the region–class label events $(R_{\tau(\theta)}(X)=r,\ Y=y)$ form a finite partition of the sample space. Consequently, for every $\theta$,
$\sum_{r\in\mathcal{R}}\sum_{y\in\mathcal{Y}}p_{r,y}(\theta)=1,\qquad\sum_{r\in\mathcal{R}}\sum_{y\in\mathcal{Y}}\widehat{p}^{\,\mathrm{audit}}_{r,y}(\theta)=1.$
Moreover, the class label marginal probabilities are fixed by the deployment data-generating process:
$\sum_{r\in\mathcal{R}}p_{r,y}(\theta)=\mathbb{P}(Y=y\mid\mathcal{D}_{\mathrm{cal}}),\qquad\sum_{r\in\mathcal{R}}\widehat{p}^{\,\mathrm{audit}}_{r,y}(\theta)=\widehat{\mathbb{P}}^{\,\mathrm{audit}}(Y=y).$
Varying $\theta$ reallocates mass across regions within each label. These conservation laws are the source of geometric coupling and explain why sweeping $\theta$ produces a constrained set of rate vectors rather than filling an ambient volume.
(B) Policy-specific indicators.
Fix a deployment policy $\pi$. Any operational quantity of interest can be written as a Bernoulli indicator
$I_{\ell}:=g_{\ell}(R_{\tau(\theta)}(X),Y;\pi)\in\{0,1\},\qquad p_{\ell}(\theta):=\mathbb{P}\!\big(I_{\ell}=1\mid\mathcal{D}_{\mathrm{cal}}\big).$
These operational targets are auditable event rates that summarize deployment behavior (e.g., abstention, decisiveness, decisive error exposure, or class-conditional region mass). For example, the abstention indicator is $g_{\mathrm{abs}}(r,y;\pi)=\mathbf{1}\{\pi(r)=\varnothing\}$, and a decisive error indicator is $g_{\mathrm{err}}(r,y;\pi)=\mathbf{1}\{|\pi(r)|=1,\ y\notin\pi(r)\}$. The corresponding audit count is
$K^{\mathrm{audit}}_{\ell}(\theta):=\sum_{i=1}^{n_{\mathrm{audit}}}g_{\ell}(R_{\tau(\theta)}(X_{i}^{a}),Y_{i}^{a};\pi),\qquad\widehat{p}^{\,\mathrm{audit}}_{\ell}(\theta):=\frac{K^{\mathrm{audit}}_{\ell}(\theta)}{n_{\mathrm{audit}}}.$
Projection identity.
The two representations are linked by linearity: every linear KPI is obtained by summing selected region–label cells. On the audit sample,
$K^{\mathrm{audit}}_{\ell}(\theta)=\sum_{r\in\mathcal{R}}\sum_{y\in\mathcal{Y}}g_{\ell}(r,y;\pi)\,K^{\mathrm{audit}}_{r,y}(\theta),\qquad\widehat{p}^{\,\mathrm{audit}}_{\ell}(\theta)=\frac{K^{\mathrm{audit}}_{\ell}(\theta)}{n_{\mathrm{audit}}}.$ (1)
Taking expectation yields the corresponding identity for latent rates:
$p_{\ell}(\theta)=\sum_{r\in\mathcal{R}}\sum_{y\in\mathcal{Y}}g_{\ell}(r,y;\pi)\,p_{r,y}(\theta).$
Thus the region–label table is a reusable sufficient audit summary: audit it once, then compute policy-level KPIs by projection. The same structure extends to monotone transformations of linear KPIs, since the region–class label table remains the sufficient summary.
Worked example: coverage as a projection of the region–label table.
Consider $\mathcal{Y}=\{0,1\}$ and the set-inclusion policy $\pi_{\mathrm{SI}}$. For fixed thresholds, each input falls into one of four regions $\mathcal{R}=\{r_{10},r_{11},r_{01},r_{00}\}$, and the primitive auditing object is $(R_{\tau}(X),Y)\in\mathcal{R}\times\mathcal{Y}$. The corresponding region–label table is
$P(\theta):\begin{array}[]{c|ccc}&y=0&y=1&\hat{C}_{\pi_{\mathrm{SI}}}(X)\\ \hline\cr r_{10}=(1,0)&\mathbf{p_{10,0}(\theta)}&p_{10,1}(\theta)&\{0\}\\ r_{11}=(1,1)&\mathbf{p_{11,0}(\theta)}&\mathbf{p_{11,1}(\theta)}&\{0,1\}\\ r_{01}=(0,1)&p_{01,0}(\theta)&\mathbf{p_{01,1}(\theta)}&\{1\}\\ r_{00}=(0,0)&p_{00,0}(\theta)&p_{00,1}(\theta)&\varnothing\end{array}\qquad\text{with}\qquad\sum_{r\in\mathcal{R}}\sum_{y\in\mathcal{Y}}p_{r,y}(\theta)=1.$
Under $\pi_{\mathrm{SI}}$, the event $\{Y\in\hat{C}_{\pi_{\mathrm{SI}}}(X)\}$ corresponds to the four bold-indicated cells $(r_{10},0)$, $(r_{11},0)$, $(r_{11},1)$, and $(r_{01},1)$, hence
$p_{\mathrm{cov}}(\theta)=p_{10,0}(\theta)+p_{11,0}(\theta)+p_{11,1}(\theta)+p_{01,1}(\theta)=1-\big(p_{10,1}(\theta)+p_{01,0}(\theta)+p_{00,0}(\theta)+p_{00,1}(\theta)\big).$
Note that the calibration choices $(\alpha_{0}^{\star},\alpha_{1}^{\star})$ determine the class-conditional miscoverage primitives—namely the joint rates $p_{10,1}(\theta)+p_{00,1}(\theta)$ for $Y=1$ and $p_{01,0}(\theta)+p_{00,0}(\theta)$ for Y=0Y=0—while the decomposition of these totals into wrong-singleton errors versus abstentions (including the overall empty-set rate $p_{00,0}(\theta)+p_{00,1}(\theta)$) is dictated by the deployment distribution. The above example makes explicit both (i) the projection form (coverage is a sum of cells) and (ii) the conservation constraint: once thresholds are fixed, coverage is determined by how mass is redistributed across regions.
The same “sum selected cells” structure applies to abstention/deferral, decisiveness, and decisive error exposure under any fixed policy. In this sense, these operational quantities are all derived from a single auditable summary, from which many KPIs can be computed by projection while keeping the conservation constraints behind the trade-offs explicit.
3 Operational quantities as first-class objects
Deployment behavior is summarized by operational rates induced by a calibrated threshold geometry together with a fixed deployment policy $\pi$. Consistent with Section 2, we keep three layers distinct: (i) geometry, given by thresholds and the induced region map $R_{\tau(\theta)}$; (ii) policy, given by a deterministic projection $\pi:\mathcal{R}\to 2^{\mathcal{Y}}$; and (iii) rates, which quantify how often auditable region–label events occur under deployment.
The goal of this section is to make operational quantities first-class: to specify which rate statements are certifiable, which data split is required, and how trade-offs are explored without committing to a scalar objective. Our perspective is explicitly predictive: we seek conservative, distribution-free statements about future realized operational behavior over finite windows, conditional on the one-time calibration draw. The intended output is an auditable trade-off map $\theta\mapsto\mathbf{r}(\theta)$ linking calibration choices to operational profiles, together with an efficiency boundary. Here $\theta\in\Theta$ denotes a calibration setting, i.e., an index for the deployed thresholds $\tau(\theta)$ and induced region map $R_{\tau(\theta)}$.
3.1 Calibrate–and–Audit
Beyond marginal coverage, distribution-free finite-sample statements require an independent audit split. Coverage is special because its split-conformal guarantee is rank-based. For general region–class label rates, reusing $\mathcal{D}_{\mathrm{cal}}$ couples region–class label indicators to the random threshold and breaks the exchangeable Bernoulli model (Geisser, 1993; Casella and Berger, 2002) needed for exact predictive envelopes; therefore certified envelopes require an independent $\mathcal{D}_{\mathrm{audit}}$ (Appendix D).
Using the audit split defined in Section 2, we keep thresholds (and thus $R_{\tau(\theta)}$) fixed from $\mathcal{D}_{\mathrm{cal}}$ and compute all operational rate estimates and predictive envelopes on $\mathcal{D}_{\mathrm{audit}}$ only. As with the train/calibrate separation, once the rule depends on data, certification of downstream behavior requires fresh exchangeable evaluation data.
When a separate $\mathcal{D}_{\mathrm{audit}}$ cannot be afforded, we provide a conservative procedure that aims to decouple threshold selection from rate estimation by evaluating each calibration point under a leave-one-out (LOO) proxy threshold that does not use that point. This proxy is intended for feasibility exploration rather than certification. In simulations, this procedure tracks the certified two-split reference closely; see Appendix D.
3.2 Certified predictive envelopes for future windows
Section 2 defined the auditable primitive: the joint process $(R_{\tau(\theta)}(X),Y)$ under fixed thresholds, together with KPI indicators obtained as linear projections of the corresponding region–class label table, with coverage as a worked example. We now use this object to construct predictive envelopes for future operational behavior. Under Calibrate–and–Audit, conditional on $\mathcal{D}_{\mathrm{cal}}$ the deployed rule (hence $R_{\tau(\theta)}$ and $\pi$) is fixed, and the audit points are exchangeable with future deployment points. Therefore, for any fixed indicator $I_{\ell}$,
$K^{\mathrm{audit}}_{\ell}(\theta)\mid\mathcal{D}_{\mathrm{cal}}\sim\mathrm{Binomial}\!\big(n_{\mathrm{audit}},p_{\ell}(\theta)\big),$
and for a future operational window of size $m$ with draws $(X^{\prime}_{j},Y^{\prime}_{j})$, the future realized count
$K^{m}_{\ell}(\theta):=\sum_{j=1}^{m}g_{\ell}(R_{\tau(\theta)}(X^{\prime}_{j}),Y^{\prime}_{j};\pi)\ \Big|\ \mathcal{D}_{\mathrm{cal}}\sim\mathrm{Binomial}\!\big(m,p_{\ell}(\theta)\big).$
These conditional Binomial laws (Johnson et al., 1997) induce distribution-free finite-sample predictive envelopes for $K^{m}_{\ell}(\theta)$, and hence for the realized window rate $\widehat{R}_{\ell,m}=K^{m}_{\ell}(\theta)/m$, based on the audit count $K^{\mathrm{audit}}_{\ell}(\theta)$ observed on $\mathcal{D}_{\mathrm{audit}}$.
The Beta–Binomial form is not a prior assumption on $p_{\ell}(\theta)$, but rather a closed-form expression for the predictive distribution of $K^{m}_{\ell}(\theta)$ induced by the two-stage design: conditional on $\mathcal{D}_{\mathrm{cal}}$, the audit indicators and future-window indicators are i.i.d. Bernoulli trials with a common (unknown) success probability $p_{\ell}(\theta)$. Observing the audit count $K^{\mathrm{audit}}_{\ell}(\theta)$ therefore determines exact finite-sample predictive tails for $K^{m}_{\ell}(\theta)$ under exchangeability, which can be written in Beta–Binomial form.
As a conservative alternative, the same audit rate $\widehat{p}^{\,\mathrm{audit}}_{\ell}(\theta)$ admits Hoeffding-type deviation inequalities (Hoeffding, 1963) for the latent rate $p_{\ell}(\theta)$. These yield worst-case planning envelopes for future window rates; we record them in Appendix D.
3.3 Rate vectors, attainable operational sets, and Pareto filtering
Certified envelopes (Section 3.2) apply to any fixed region–label event rate at any fixed calibration setting, and to any linear projection thereof. Deployment planning is comparative: stakeholders want to understand how operational behavior varies across calibration settings, which combinations of rates are feasible for a given scoring model and coverage semantics, and where structural bottlenecks lie. This motivates treating the rate vector (Miettinen, 1999) as the decision interface.
Let $\Theta$ denote a family of calibration settings (e.g., an $\alpha$-grid or a class-conditional threshold family), and write $R_{\tau(\theta)}$ for the induced region map at setting $\theta$. Fix a deployment policy $\pi$ and a list of $L$ auditable KPIs, each represented by an indicator $I_{\ell}=g_{\ell}(R_{\tau(\theta)}(X),Y;\pi)$. Define the calibration-conditional rate vector
$\mathbf{r}(\theta)=\big(r_{1}(\theta),\ldots,r_{L}(\theta)\big),\qquad r_{\ell}(\theta):=\mathbb{P}\!\big(I_{\ell}=1\mid\mathcal{D}_{\mathrm{cal}}\big).$
Each coordinate is obtained by projection from the same audited region–label table via (1).
Attainable operational set and visualization.
Sweeping $\theta$ traces the set of achievable operational profiles under the fixed policy $\pi$,
$\mathcal{V}(\Theta;\pi):=\{\mathbf{r}(\theta):\theta\in\Theta\}\subset\mathbb{R}^{L}.$
Because the region–label table obeys conservation constraints (sum-to-one and fixed label marginals), $\mathcal{V}(\Theta;\pi)$ is typically a constrained set: changing $\theta$ reallocates probability mass across regions and thus across KPIs, rather than tuning rates independently. Trade-off plots (e.g., Figure 1) display selected coordinates of $\widehat{\mathbf{r}}^{\,\mathrm{audit}}(\theta)$, optionally annotated with finite-window predictive envelopes computed from the projected audit counts.
Orientation and Pareto filtering
To summarize negotiation-relevant regimes without committing to a scalar objective, we apply an oriented Pareto filter. Concretely, the analyst chooses, for each KPI, whether it is desirable to be large or small (e.g., lower decisive error exposure is better; higher purity is better). This is an orientation choice, not a pricing model. Formally, fix an orientation vector $s\in\{-1,+1\}^{L}$ and compare points using $s\odot\mathbf{r}(\theta)$. A setting $\theta$ is nondominated if there is no $\theta^{\prime}$ such that
$s_{\ell}\,r_{\ell}(\theta^{\prime})\geq s_{\ell}\,r_{\ell}(\theta)\ \ \text{for all }\ell,\qquad\text{and}\qquad s_{\ell_{0}}\,r_{\ell_{0}}(\theta^{\prime})>s_{\ell_{0}}\,r_{\ell_{0}}(\theta)\ \ \text{for some }\ell_{0}.$
The nondominated set is the negotiation front: any later decision rule that imposes hard constraints or introduces scalar weights must select an operating point on (or near) this front. Moreover, if a finalized cost or utility functional is monotone with respect to the chosen orientation, any optimum must be attained on the nondominated front (Miettinen, 1999).
The resulting trade-off map answers a planning question: within the permissible calibration family $\Theta$, what operational behaviors are attainable under the fixed policy $\pi$, and which of those behaviors survive as negotiation-relevant regimes? Predictive envelopes then attach finite-window uncertainty to the frontier points, supporting robust regime selection under operational constraints, or can be included to shape the pareto frontier directly.
3.4 SSBC as a calibration navigation coordinate
A trade-off map is most useful in practice when it admits a user-facing means of navigation tied to an agreed-upon semantic constraint. In this work we use coverage as that semantic anchor. Coverage is special among the KPIs considered above: under exchangeability, split conformal admits an exact finite-sample rank/Beta characterization of the calibration-conditional coverage (Marques, 2025; Angelopoulos and Bates, 2023). SSBC inverts this law to translate a semantic request $(\alpha^{\star},\delta)$ into a concrete discrete choice on the conformal calibration grid (Appendices B and C).
Split conformal calibration selects thresholds on a finite grid of order statistics; equivalently, it can be indexed by the discrete miscoverage level $\alpha_{\mathrm{grid}}=u/(n_{\mathrm{cal}}+1)$ with $u\in\{1,\ldots,n_{\mathrm{cal}}\}$ (or by the equivalent order-statistic index $k=n_{\mathrm{cal}}+1-u$; see Section 2.2). SSBC maps $(\alpha^{\star},\delta)$ to a single admissible index $u^{\star}(\alpha^{\star},\delta)$ by selecting the least conservative grid point whose induced calibration-conditional coverage satisfies the PAC-style tail requirement
$\mathbb{P}_{\mathcal{D}_{\mathrm{cal}}}\!\left(\mathbb{P}\!\big(Y\in\hat{C}(X)\mid\mathcal{D}_{\mathrm{cal}}\big)\geq 1-\alpha^{\star}\right)\geq 1-\delta.$
Concretely, SSBC inverts the exact rank/Beta law and returns the boundary grid point that satisfies the $\delta$-tail constraint with minimal conservatism.
This reduction is important for the geometry studied in Section 4: although a user specifies two parameters $(\alpha^{\star},\delta)$, SSBC returns a single discrete index $u^{\star}$, hence a single threshold choice, Appendix C. That choice fixes the region map $R_{\tau(\theta)}$ and therefore fixes all KPIs in $\mathbf{r}(\theta)$.
In pooled split conformal there is one such index, yielding a one-parameter navigation along the conformal grid. In class-conditional calibration, SSBC is applied separately within each class. In the binary case this yields two indices
$(u_{0}^{\star},u_{1}^{\star})=\big(u^{\star}(\alpha_{0}^{\star},\delta_{0}),\,u^{\star}(\alpha_{1}^{\star},\delta_{1})\big),$
which determine $(\tau_{0},\tau_{1})$ and hence the calibration setting $\theta$. Thus, while the user-facing specification appears four-dimensional $(\alpha_{0}^{\star},\delta_{0},\alpha_{1}^{\star},\delta_{1})$, conformal calibration collapses this into a two-dimensional navigation coordinate on the calibration grid.
Implication for attainable sets.
Because $\theta$ is determined by these discrete indices, the calibration family $\Theta$ admits a low-dimensional parameterization even before accounting for the conservation constraints in the region–class label table. Section  4 then makes the resulting coupling explicit in the probability-normalized binary setting and explains the sharp regime boundaries observed when sweeping these navigation indices.
4 Consequences of a fixed conformal partition in the binary case
This section examines the geometry induced by a fixed calibrated conformal interface. Given a trained scoring function and a realized calibration draw, calibration selects thresholds $\tau$ and therefore a fixed region process $R_{\tau}(X)$ (Section 2.3). All deployment-facing quantities in this paper are determined by the joint region–class label structure $(R_{\tau}(X),Y)$ (Section 3). In the binary case, this viewpoint has three practical implications. First, feasibility is coupled: varying $\tau$ reallocates probability mass across finitely many region types rather than tuning rates independently. Second, region-based policies depend on within-region class label composition. Third, once actions are tied to regions, coherence with a cost or utility model imposes compatibility constraints that depend on the region–class label table (Appendix F). By coherence we mean that, for each region, the chosen action is optimal among the admissible actions under the given cost structure, so that the resulting expected cost is as small as possible under the corresponding region-conditional label distribution. We make these points explicit under probability-normalized scores.
4.1 Binary conformal geometry and regime boundaries
Let $\mathcal{Y}=\{0,1\}$ and fix class-conditional thresholds $\tau=(\tau_{0},\tau_{1})$. The deployed interface is the region label
$R_{\tau}(x):=\big(\mathbf{1}\{s(x,0)\leq\tau_{0}\},\ \mathbf{1}\{s(x,1)\leq\tau_{1}\}\big)\in\{0,1\}^{2},$
with outcomes $10,11,01,00$ (singleton-0, hedge, singleton-$1$, abstention under $\pi_{\mathrm{SI}}$).
Probability-normalized scores induce a regime boundary.
For probability-normalized scores $s(x,y)=1-P(y\mid x)$ we have $s(x,0)+s(x,1)=1$. Hence a point cannot satisfy both class thresholds unless $\tau_{0}+\tau_{1}\geq 1$, and it cannot violate both unless $\tau_{0}+\tau_{1}\leq 1$. Consequently:
  • Hedging regime ($\tau_{0}+\tau_{1}>1$): $11$ may occur and $00$ cannot (singletons + hedges; no abstention).
  • Rejection regime ($\tau_{0}+\tau_{1}<1$): $00$ may occur and $11$ cannot (singletons + abstention; no hedges).
  • Boundary ($\tau_{0}+\tau_{1}=1$): only $10$ and $01$ occur; under $\pi_{\mathrm{SI}}$ outputs are always singletons.
Crossing $\tau_{0}+\tau_{1}=1$ therefore changes which outcome types can appear. This is a deterministic constraint induced by the score normalization, not an estimation artifact. Note that when $\tau_{0}=\tau_{1}=1/2$, the associated classification rule is the argmax rule and coverage equals the classifier’s accuracy.
Cross-threshold coupling within regimes.
Within the hedging regime ($\tau_{0}+\tau_{1}>1$), the diagonal support is partitioned into three contiguous intervals. Parameterize by $u=s(x,0)$ (so s(x,1)=1us(x,1)=1-u):
$u\in[0,1-\tau_{1})\ \Rightarrow\ R_{\tau}(x)=10,\qquad u\in[1-\tau_{1},\tau_{0}]\ \Rightarrow\ R_{\tau}(x)=11,\qquad u\in(\tau_{0},1]\ \Rightarrow\ R_{\tau}(x)=01.$
Thus region boundaries are governed by opposing thresholds: changing $\tau_{1}$ moves the boundary that controls the mass of $10$, while changing $\tau_{0}$ moves the boundary that controls the mass of $01$. Thresholds therefore act as mass-reallocation boundaries, not independent per-class knobs.
Sweeping calibration settings moves $\tau$ and reallocates mass across regions. When score support is low-dimensional, like in the binary case with probability-normalized scores, small changes in $\tau$ can trigger support discontinuities, and within regimes operational coordinates are coupled. Under probability-normalized scores, swapping $(\tau_{0},\tau_{1})$ exchanges the roles of regions $11$ and $00$. If the deployment policy treats hedging and abstention identically, then this swap does not change action-level behavior, even though coverage changes.
4.2 Region–label tables, piecewise-constant posteriors, and coherent action
Fix $\tau$ and treat the region label $R_{\tau}(X)\in\{0,1\}^{2}$ as the only deployed observable. The joint region–label probabilities
$p_{r,y}:=\mathbb{P}\!\left(R_{\tau}(X)=r,\ Y=y\mid\mathcal{D}_{\mathrm{cal}}\right),\qquad r\in\{0,1\}^{2},\ y\in\{0,1\},$
fully characterize the information available to any region-associated policy. Write $p_{r}:=p_{r,0}+p_{r,1}$ for region mass. Because the interface takes finitely many values, the region-conditioned label composition is constant within each region; downstream evaluation and decision-making therefore reduce to functions of $(p_{r,0},p_{r,1})$, aggregated across regions using the same table $\{p_{r,y}\}$.
Cost-coherence for region-based action.
Let $\mathcal{A}$ be an action set and let $L(a,y)$ be the cost of taking action $a\in\mathcal{A}$ when $Y=y$. A region-based policy $\tilde{\pi}:\{0,1\}^{2}\to\mathcal{A}$ induces calibration-conditional expected cost
$\mathcal{R}(\tilde{\pi}):=\sum_{r\in\{0,1\}^{2}}\sum_{y\in\{0,1\}}L(\tilde{\pi}(r),y)\,p_{r,y}.$
We say $\tilde{\pi}$ is cost-coherent for the deployed interface if, for every region with $p_{r}>0$, it minimizes conditional risk within that region. In words, cost coherence means choosing, in each region, an action that minimizes expected cost under the corresponding region-conditional label distribution:
$\tilde{\pi}(r)\in\arg\min_{a\in\mathcal{A}}\frac{\sum_{y\in\{0,1\}}L(a,y)\,p_{r,y}}{p_{r}}\qquad\text{for all $r$ with $p_{r}>0$.}$
[…]
Chow-style Acceptance/Rejection
Following Chow (1970), consider actions $\mathcal{A}=\{0,1,\mathrm{rej}\}$ with false positive cost $c_{10}>0$, false negative cost $c_{01}>0$, rejection cost $c_{\mathrm{rej}}\geq 0$, and zero cost for correct commitments. For each region $r$, define $p_{r}:=p_{r,0}+p_{r,1}$. Cost-coherence is decided by comparing the three region-level alternatives:
$\text{choose }0:\quad p_{r,1}c_{01}\leq p_{r,0}c_{10}\ \text{ and }\ p_{r,1}c_{01}\leq p_{r}c_{\mathrm{rej}},$
$\text{choose }1:\quad p_{r,0}c_{10}\leq p_{r,1}c_{01}\ \text{ and }\ p_{r,0}c_{10}\leq p_{r}c_{\mathrm{rej}},$
$\text{choose rej}:\quad p_{r}c_{\mathrm{rej}}\leq p_{r,1}c_{01}\ \text{ and }\ p_{r}c_{\mathrm{rej}}\leq p_{r,0}c_{10}.$
Thus, for a fixed convention, coherence means its prescribed action satisfies the corresponding inequalities in every region (ties allowed). For example, if the policy commits to class 0 in region $10$ (singleton output $\{0\}$ under $\pi_{\mathrm{SI}}$), then coherence requires $p_{10,1}c_{01}\leq p_{10,0}c_{10}$ and $p_{10,1}c_{01}\leq(p_{10,0}+p_{10,1})c_{\mathrm{rej}}$. This also shows why the hard-wired convention “commit on singleton outputs; defer otherwise” is cost-coherent only if, in each singleton region with $p_{r}>0$, the forced commitment has lower expected cost than both rejection and the opposite commitment: a region can be labeled by a singleton set under $\pi_{\mathrm{SI}}$ yet still favor rejection (or even the opposite commitment) for plausible cost ratios.
This example illustrates the basic point behind cost-coherence: once a conformal partition is deployed, whether a downstream action convention is rational depends on the information carried by the region label together with the costs of the available actions. In particular, even if a region is labeled as a singleton under $\pi_{\mathrm{SI}}$, coherence is determined by the region–label composition $(p_{r,0},p_{r,1})$, not by the set label alone.
More generally, coherence of any fixed (interface, convention) pair can be expressed as explicit inequalities in the region–class label table $\{p_{r,y}\}$ and cost ratios such as $(c_{\mathrm{rej}}/c_{01},c_{\mathrm{rej}}/c_{10})$; see Appendix F. These inequalities restrict the range of cost ratios under which a given convention can be cost-coherent. The appendix develops this inverse view by characterizing, for a fixed convention, the corresponding coherent region of cost ratios.
5 Results: validation and deployment-facing planning
This section reports results aligned with the paper’s deployment-facing claims for a single deployed scoring model: the model is treated as fixed infrastructure and we vary only calibration settings. We focus on three questions: (i) SSBC coverage semantics under calibration randomness and finite deployment windows, (ii) predictive envelopes for deployment-facing operational rates beyond coverage, and (iii) scenario planning via trade-off maps with cost-coherence screening.
5.1 Numerical Simulations
This subsection validates two core claims: SSBC preserves the intended finite-window coverage semantics under calibration randomness, and Calibrate–and–Audit provides predictive envelopes for deployment-facing operational rates. Table 1 summarizes semantic coverage behavior, and Figure 4 compares envelope alignment between two-sample and LOO constructions.
5.1.1 Coverage: numerical realization of SSBC guarantees
This experiment instantiates the exact finite-sample laws underlying SSBC in the finite-window deployment regime, separating calibration randomness from deployment-window noise. Calibration nonconformity scores are drawn i.i.d. from a continuous, heavy-tailed reference distribution, chosen here as the absolute Cauchy distribution $|t_{\nu=1}|$.
For each calibration set size $n_{\mathrm{cal}}$, split conformal selects a threshold via a discrete miscoverage index $u\in\{1,\ldots,n_{\mathrm{cal}}\}$ with
$\alpha_{\mathrm{grid}}=\frac{u}{n_{\mathrm{cal}}+1},\qquad k=n_{\mathrm{cal}}+1-u,\qquad\tau=S_{(k)}.$
We compare three index-selection rules at target $(\alpha^{\star},\delta)=(0.10,0.10)$: nominal split conformal, a DKWM correction (Massart, 1990; Dvoretzky et al., 1956), and SSBC, Appendix C. In the latter case, we distinguish between the infinite window and finite window cases that follow a Beta and Beta-Binomial law, respectively.
Table 1 shows representative calibration sizes (full grid in Appendix C). Nominal split conformal under-controls calibration-conditional risk; DKWM enforces validity through strong conservatism. SSBC directly inverts the relevant finite-window law, yielding violation probabilities close to the target $\delta$ up to grid effects.
Table 1: Representative calibration-conditional coverage violation rates. Target miscoverage $\alpha^{\star}=0.10$, confidence $\delta=0.10$, and deployment window size $m_{\mathrm{infer}}=100$. We report three representative calibration sizes; the complete grid is in Table 7. Obs, Beta and BetaBinom are the observed, theoretical, and Beta-Binomial theoretical violation probabilities, respectively. SSBC achives near-target violation probabilities for all calibration sizes, consistent with Beta-Binomial theory.
$n_{\mathrm{cal}}$ Method $u$ $\alpha_{\mathrm{grid}}$ Obs Beta BetaBinom $\alpha_{\mathrm{cont}}$
100 None 10 0.0990 0.4075 0.4513 0.4071
100 SSBC 6 0.0594 0.0960 0.0576 0.0956
100 DKWM 1 0.0099 0.0004 0.0000 0.0004 0.0224-0.0224
200 None 20 0.0995 0.4130 0.4655 0.4124
200 SSBC 13 0.0647 0.0974 0.0320 0.0980
200 DKWM 2 0.0100 0.0000 0.0000 0.0000 +0.0135+0.0135
500 None 50 0.0998 0.4153 0.4782 0.4153
500 SSBC 34 0.0679 0.0949 0.0049 0.0955
500 DKWM 22 0.0439 0.0097 0.0000 0.0096 +0.0453+0.0453
5.1.2 Operational rate envelopes: LOO versus two-sample reference
We next validate predictive envelopes for operational rates beyond coverage, where no distribution-free rank pivot exists. The goal is to compare the certified two-sample Calibrate–and–Audit construction to a single-sample LOO surrogate intended for feasibility exploration.
Synthetic probability model.
We generate predicted probabilities and labels so the score geometry is controlled and interpretable. Each draw produces $Y\in\{0,1\}$ with $\mathbb{P}(Y=1)=p_{\mathrm{class}}$, where $p_{\mathrm{class}}\in\{0.10,0.50\}$. Conditional on $Y$, we draw $P_{1}$ via
$P_{1}\mid Y=1\sim\mathrm{Beta}(a,b),\qquad P_{1}\mid Y=0\sim\mathrm{Beta}(2,7),$
with $(a,b)\in\{(4,3),(9,3)\}$, form $(P_{0},P_{1})=(1-P_{1},P_{1})$, and use class-conditional scores $S_{y}:=1-P_{y}$.
Two-sample reference and LOO surrogate.
For each configuration, we draw independent datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ of size $N=500$. We calibrate on $\mathcal{D}_{1}$ (including SSBC-adjusted thresholds at $(\alpha,\delta)=(0.10,0.10)$ in the finite-window regime), freeze the induced rule, and evaluate operational indicators on $\mathcal{D}_{2}$, yielding exact Binomial/Beta–Binomial predictive envelopes. As a single-sample surrogate, we compute LOO operational indicators by evaluating each point under thresholds recomputed without that point, pool LOO counts, and map them to envelopes; optionally we widen intervals via controlled pessimization (Appendix D).
Figure 4 shows that LOO-based envelopes closely match the two-sample reference in both location and width across all four geometries, while inflation widens intervals without shifting centers. This supports using LOO as a practical proxy for planning when an explicit audit split is unavailable, while keeping certification tied to the two-sample design.
Figure 4: Operational rate envelopes and underlying score distributions. (A) Singleton rate and (B) singleton error for two class prevalences ($p_{\mathrm{class}}\in\{0.10,0.50\}$) and two class 1 generating distributions ($\mathrm{Beta}(4,3)$ and $\mathrm{Beta}(9,3)$), with class 0 generated by $\mathrm{Beta}(2,7)$. Red rectangles denote the two-sample Calibrate–and–Audit Beta–Binomial ($B$) predictive envelope, and the dashed vertical line is the corresponding $B$ point estimate. Blue and orange markers with horizontal intervals show leave-one-out (LOO) envelopes computed from a single calibration dataset under two inflation levels ($\mathrm{infl}=1,2$). (C1–C4) Histograms of the score geometry shown as predicted class 1 probability (equivalently, nonconformity to class 0) stratified by true class. These distributions make explicit how class prevalence and calibration geometry shape singleton mass and singleton error, and explain the regime-dependent asymmetry of the finite-sample envelopes.
These simulations support the calibration and envelope semantics used in the two real-data studies below.
5.2 Tox21: predictive envelopes under severe class imbalance
The Tox21 benchmark (Mayr et al., 2016; Huang et al., 2016) stress-tests the framework under severe class imbalance, where minority-class calibration counts can be well below $100$ (Table 2). We evaluate (i) SSBC semantic coverage behavior under small class-conditional calibration sizes and (ii) finite-window operational envelopes in a practical setting. Table 3 summarizes aggregated semantic behavior; endpoint-level operational envelope comparisons appear in Table 4 and Appendix H.
Tasks and modeling infrastructure.
Tox21 comprises twelve binary toxicity endpoints (NR and SR pathways). Each endpoint is treated as an independent binary classification problem. Molecules are represented using RDKit (Landrum, 2025) 2D descriptors augmented with Morgan fingerprints (Rogers and Hahn, 2010) (radius $2$, $128$ bits), and classification is performed using CatBoost (Prokhorenkova et al., 2018) with fixed hyperparameters. For each endpoint, data are split randomly into training (50%), calibration (25%), and audit (25%), and results are averaged over $100$ splits. Conformal calibration is class-conditional and we compare: (i) standard split conformal, (ii) DKWM correction, and (iii) SSBC, all at $(\alpha,\delta)=(0.10,0.10)$.
Table 2: Tox21 dataset composition and effective class-conditional calibration sizes under the experimental protocol.
Endpoint Total Samples Positive Rate Calib. Positives Calib. Negatives
NR-AR 7265 4.3% 77 1739
NR-AR-LBD 6758 3.5% 59 1630
NR-AhR 6549 11.7% 192 1445
NR-Aromatase 5821 5.2% 75 1380
NR-ER 6193 12.8% 198 1350
NR-ER-LBD 6955 5.0% 87 1651
NR-PPAR-$\gamma$ 6450 2.9% 46 1566
SR-ARE 5832 16.2% 235 1223
SR-ATAD5 7072 3.7% 66 1702
SR-HSE 6467 5.8% 93 1523
SR-MMP 5810 15.8% 229 1223
SR-p53 6774 6.2% 105 1587
5.2.1 Coverage semantics under small class-conditional calibration sizes
Table 3 aggregates empirical coverage, violation rates, and prediction-set statistics over all endpoints and splits. Nominal split conformal exhibits elevated violation probability in this conditional small-$n$ regime. DKWM enforces validity through conservatism, inflating set sizes and reducing singleton frequency. SSBC substantially reduces violation probability while retaining more decisiveness than DKWM.
Table 3: Aggregated coverage and prediction-set statistics on Tox21. Results are averaged over twelve endpoints and 100 random splits. Violation rate is the empirical probability that coverage falls below the target level. Both $\alpha$ and $\delta$ are set to 0.10.
Method Mean Coverage Violation Rate Avg. Set Size Singleton Rate
Standard Split Conformal 0.917 0.305 1.41 0.52
DKWM Correction 0.986 0.005 1.78 0.22
SSBC 0.951 0.068 1.54 0.40
5.2.2 Operational envelopes as predictive summaries
With coverage risk fixed, we evaluate predictive envelopes for operational quantities over a finite deployment window. Operational KPIs are region–label indicators (Section 2.4); in the binary case, singleton mass, doublet mass, and wrong-singleton mass provide a compact KPI set tied directly to decisiveness and decisive error exposure.
We compare the two-sample Calibrate–and–Audit reference to the single-sample LOO surrogate (Appendix D). Table 4 provides a representative endpoint-level breakdown; an additional endpoint is reported in Appendix H (Table 8). LOO point estimates track two-sample observed rates and intervals are comparable (often slightly wider), consistent with LOO primarily affecting predictive dispersion rather than centering.
Table 4: SR-MMP endpoint. Operational performance summary for the SR-MMP endpoint. All reported quantities are joint probabilities normalized by the total test-set size. Rows report the singleton and doublet joint event rates P(Y=c,|S|=1)P(Y=c,\,|S|=1) and P(Y=c,|S|=2)P(Y=c,\,|S|=2), together with the wrong-singleton rate $P(Y=c,\,|S|=1,\,\hat{y}\neq Y)$, by true class. “point estimate $\mathcal{D}_{1}$” and “LOO 95% PI $\mathcal{D}_{1}$” denote leave-one-out point estimates and their beta–binomial predictive intervals computed on the calibration data. “Observed $\mathcal{D}_{2}$” reports empirical rates on an independent hold-out set. “$B$ 95% PI $\mathcal{D}_{2}$” gives the beta–binomial predictive intervals for the corresponding $\mathcal{D}_{2}$ quantities, accounting for calibration uncertainty and finite test size.
Operational quantity Class point estimate $\mathcal{D}_{1}$ LOO 95% PI $\mathcal{D}_{1}$ Observed $\mathcal{D}_{2}$ BB 95% PI $\mathcal{D}_{2}$
Singleton rate Class 0 $0.662$ $[0.604,0.718]$ $0.651$ $[0.615,0.686]$
Class 1 $0.130$ $[0.092,0.173]$ $0.117$ $[0.094,0.142]$
Doublet rate Class 0 $0.163$ $[0.121,0.211]$ $0.201$ $[0.172,0.231]$
Class 1 $0.045$ $[0.023,0.074]$ $0.031$ $[0.020,0.046]$
Wrong-singleton rate Class 0 $0.073$ $[0.044,0.107]$ $0.070$ $[0.052,0.090]$
Class 1 $0.012$ $[0.002,0.030]$ $0.011$ $[0.005,0.020]$
5.3 R3: scenario planning on aqueous solubility
We now demonstrate the planning use-case: exploring attainable deployment-facing trade-offs across calibration settings without committing to a single scalar objective. We focus on aqueous solubility in drug development using AquaSolDB (Sorkun et al., 2019). The predictive model is treated as fixed infrastructure; variation is entirely through the calibration layer (SSBC), the induced conformal partition, and the downstream interpretation of the resulting finite interface.
Methodological setup (documented in Appendix I).
This experiment is designed to separate (i) generalization across chemical space from (ii) exchangeable evaluation for conformal and finite-window forecasting. We therefore use a random train/calibration/test split, with the calibration and test sets serving as the exchangeable evaluation design used throughout the solubility experiments (Appendix I, dataset/partitioning/model details). Although the base classifier is trained on a three-class discretization of $\log S$ (Insoluble / Moderate / Soluble), deployment-facing planning is posed as a binary campaign objective by merging Moderate and Soluble into a single Soluble class and calibrating prediction sets in $\{\mathrm{Ins},\mathrm{Sol}\}$ (Appendix I, calibration and operationalization).
For the operational planning demonstration, we simulate a campaign focused on lipophilic compounds. Concretely, we take the focal deployment scenario to be $\{\mathrm{MolLogP}>3.5\}$ and compute SSBC thresholds, together with all derived predictive envelopes, using the calibration sample restricted to this subpopulation (Appendix I, Section I.6).
Operating points and audited outcomes.
For each SSBC setting, calibration fixes thresholds and therefore a discrete binary conformal interface ($\{\mathrm{Sol}\}, singleton-$$\{\mathrm{Ins}\}$, or doublet $\{\mathrm{Sol},\mathrm{Ins}\}$). We interpret singletons as decisive commitments and the doublet as deferral, and summarize each setting by the joint outcome allocation $P(Y=y,C=c)$ over $y\in\{\mathrm{Sol},\mathrm{Ins}\}$ and $c\in\{\{\mathrm{Sol}\},\{\mathrm{Ins}\},\{\mathrm{Sol},\mathrm{Ins}\}\}$ (Table 5).
Table 5: Operational outcome categories induced by binary prediction sets.
True label $Y$ $C(X)=\{\mathrm{Sol}\}$ $C(X)=\{\mathrm{Ins}\}$ $C(X)=\{\mathrm{Ins},\mathrm{Sol}\}$
$\mathrm{Sol}$ TP (correct Sol singleton) FN (loss) HP (hedged Sol)
$\mathrm{Ins}$ FP (waste) TN (correct Ins singleton) HN (hedged Ins)
Joint rates are the auditable primitive: they form a complete mass allocation over the set–label outcome space and map directly to expected finite-window counts (e.g., $m\times P(Y=y,C=c)$ for a window of size $m$). Because the conformal grid is discrete and behavior is determined by threshold geometry, many nominally distinct $(\alpha_{0},\delta_{0},\alpha_{1},\delta_{1})$ settings collapse to identical realized interfaces and joint allocations; we therefore deduplicate settings and retain the least conservative representative among equivalence classes (Appendix I).
Trade-off map and cost-coherence analysis.
Figure 5 embodies the main results of this section. The left panel plots the attainable set in a KPI plane spanned by two soluble-class joint rates: irreversible exclusion $P(Y=\mathrm{Sol},C=\{\mathrm{Ins}\})$ and deferral burden $P(Y=\mathrm{Sol},C=\{\mathrm{Sol},\mathrm{Ins}\})$, with color encoding the decisive-correct soluble mass $P(Y=\mathrm{Sol},C=\{\mathrm{Sol}\})$. We highlight nondominated regimes using an orientation-only Pareto filter over these three soluble-class quantities, yielding a shortlist of negotiation-relevant operating regimes.
We also take the inverse pricing view discussed in Section 4.2. For the fixed downstream action convention $\{\mathrm{Sol}\}\!\to\!\mathrm{Sol}$, $\{\mathrm{Ins}\}\!\to\!\mathrm{Ins}$, $\{\mathrm{Sol},\mathrm{Ins}\}\!\to\!\mathrm{rej}$, the right panel characterizes the relative cost ratios $\lambda=c_{01}/c_{10}$ and $\rho=c_{\mathrm{rej}}/c_{10}$ for which this convention is cost-coherent. The result shows that points on the Pareto front are coherent only over a limited region of cost-ratio space: if the cost of irreversible exclusion is too high relative to the cost of deferral, or if the cost of deferral is too high relative to the cost of irreversible exclusion, then the fixed action convention is no longer cost-minimizing in those regions.
Additional results and documentation (appendix).
Appendix I provides the supporting methodological record and supplementary empirical summaries used to interpret Figure 5: (i) dataset and label construction, random partitioning, features, and model protocol; (ii) scenario-conditional (tribe-based) calibration restriction and its interpretation; and (iii) diagnostics on the functional roles of $(\alpha_{0},\delta_{0},\alpha_{1},\delta_{1})$ along the Pareto front, together with a discussion of how $\alpha$- and $\delta$-variation act as coarse versus fine adjustments under finite-sample SSBC geometry (Appendix I, Sections I.9I.10).
Figure 5: Solubility scenario planning via conformal operating regimes and cost-coherence. Left: attainable joint-rate map induced by sweeping SSBC calibration settings (after deduplication). Axes report joint rates on soluble compounds: irreversible exclusion $P(Y=\mathrm{Sol},C=\{\mathrm{Insol}\})$ (x-axis) and deferral burden $P(Y=\mathrm{Sol},C=\{\mathrm{Sol},\mathrm{Insol}\})$ (y-axis). Color encodes the joint decisive-correct rate $P(Y=\mathrm{Sol},C=\{\mathrm{Sol}\})$. Red rings mark Pareto-optimal operating points; thick rings mark unique KPI regimes, labeled by regime_id:multiplicity. Right: cost-coherence landscape for the fixed downstream convention $\{\mathrm{Sol}\}\!\to\!\mathrm{Sol},\ \{\mathrm{Insol}\}\!\to\!\mathrm{Insol},\ \{\mathrm{Sol},\mathrm{Insol}\}\!\to\!\mathrm{rej}$, parameterized by cost ratios $\lambda=c_{01}/c_{10}$ (irreversible loss / downstream waste) and $\rho=c_{\mathrm{rej}}/c_{10}$ (deferral / downstream waste). The heatmap shows the fraction of Pareto regimes for which this convention is cost-coherent given only the conformal output (via within-region label composition). Blue and red outlines give the union and intersection of feasible $(\lambda,\rho)$ regions across the Pareto front; thin outlines show representative wedges for the unique KPI regimes.
6 Discussion
In a deployed conformal system, the object of interest is not only marginal coverage but the realized behavior of a fixed calibrated rule over a finite operational window. In that setting, coverage is special: because it depends only on the rank of the future true-label score among calibration scores, its calibration-conditional variability admits an exact distribution-free finite-sample characterization. Most other deployment quantities, such as commitment, abstention, and error exposure, do not share this pivot structure; they depend on score geometry and threshold interactions, and therefore require explicit auditing and uncertainty quantification.
One consequence is that calibration should be viewed as selecting a fixed interface rather than merely a scalar guarantee. Once thresholds are chosen, the induced conformal partition constrains which operational profiles are attainable. Not all combinations of commitment, deferral, and decisive error exposure can be reached by tuning a single calibration parameter, because changing thresholds reallocates mass across a finite set of region types rather than adjusting each KPI independently. The regime structure seen in the trade-off curves is therefore a property of the induced geometry, not an artifact of a particular synthetic example.
This viewpoint also clarifies the roles of SSBC and Calibrate–and–Audit. SSBC addresses the mismatch between a user’s requested coverage level and what nominal split conformal delivers when calibration sets are small, by selecting the calibration index through inversion of the exact Beta or Beta–Binomial laws. The audit step addresses a different issue: for quantities beyond coverage, uncertainty must be propagated through the realized region–class label table rather than through a rank pivot. In practice, this produces a way to compare calibration settings, quantify finite-window uncertainty for several KPIs at once, and examine cost-coherence of downstream conventions without retraining the base model.
The same geometry also explains why cost-aware operation can fail. Whether a downstream action convention is coherent depends on the within-region label frequencies, not on the conformal set label by itself. A singleton region may still support rejection, or even the opposite commitment, for plausible cost ratios. In that sense, the audit table is not only a device for uncertainty quantification; it also makes explicit which downstream actions are compatible with the information content of the deployed interface.
Several limitations delimit the scope of these results. We focused on binary classification for clarity; extending the same analysis to multi-class outputs, structured predictions, or more complex abstention policies is an important next step. Audit-based predictive envelopes can also become noisy when region masses are sparse, so rare regions may require pooling, hierarchical regularization, or additional data collection. SSBC remains tied to the discrete split-conformal grid, so very strict $(\alpha^{\star},\delta)$ targets may be unattainable at small $n_{\mathrm{cal}}$. More broadly, all of the distribution-free statements rely on exchangeability; dataset shift or model drift (Tibshirani et al., 2019; Fannjiang et al., 2022) breaks the exact finite-sample laws and must be addressed through monitoring, recalibration, or shift-aware variants (Vovk, 2015).
Taken together, these results point to a useful division of labor. Coverage can be calibrated exactly in the split-conformal setting, but broader questions about how the deployed system behaves must be answered by auditing the realized interface. That distinction matters whenever deployment decisions depend on operational trade-offs rather than on coverage alone.
7 Conclusion
The main message of the paper is that a deployed conformal predictor should be treated as a fixed operational interface, not only as a coverage certificate. In the setting studied here, SSBC calibrates the coverage target at finite sample, while Calibrate–and–Audit is used to evaluate the resulting rule in terms that matter for use: commitment, abstention, error exposure, and decision coherence.
The present analysis is intentionally narrow. It relies on exchangeability and focuses on binary classification, so extending the same operational viewpoint to shift-robust procedures, richer output spaces, and more flexible downstream policies is the natural next step.
Acknowledgements
This work was supported by the U.S. Department of Energy, Office of Science, under Contract No. DE-AC02-05CH11231. Support also came in part from the Berkeley Synchrotron Infrared Structural Biology Program (BSISB), funded by the Biological and Environmental Research Division of the U.S. Department of Energy Office of Science, and from Laboratory Directed Research and Development (LDRD) funding. Additional support came from the U.S. Department of Energy, Office of Science, Scientific Discovery through Advanced Computing (SciDAC) program FORUM-AI for first-principles calculations and AI model development.
Appendix A Appendix Outline & Notation
This appendix serves as a notation guide for the technical appendices that follow. The remaining material is organized as follows:
  • SSBC and coverage pivots: derivations and calibration-conditional guarantees for grid selection, finite-sample behavior, and pivot-based uncertainty control.
  • Operational rates and LOO justification: construction of deployment-facing rate envelopes, together with single-sample leave-one-out surrogates and their assumptions.
  • Geometry and binary partitions: structural analysis for region definitions and indicator decompositions used in operational auditing.
  • Empirical case studies: additional experimental details and results for Tox21 and Solubility.
The notation below is shared across split conformal calibration, the Small-Sample Beta Correction (SSBC), and finite-window prediction of operational rates. The analysis mixes three layers: (i) discrete split-conformal thresholding via order statistics on calibration scores; (ii) SSBC grid selection and calibration-conditional coverage semantics; and (iii) predictive envelopes for deployment-facing operational rates (Calibrate-and-Audit and a single-sample leave-one-out (LOO) surrogate). To avoid collisions, we reserve $(k,u)$ exclusively for split conformal thresholding. Pooled LOO indicator counts use $N_{\mathrm{loo}}$.
Table 6: Key notation used across calibration, SSBC, and operational analysis.
Symbol Meaning Scope / Remarks
Data splits and window sizes
$\mathcal{D}_{\mathrm{cal}}$ Calibration dataset Exchangeable sample used to set conformal thresholds; size $n_{\mathrm{cal}}$.
$\mathcal{D}_{\mathrm{audit}}$ Audit / evaluation dataset Exchangeable sample used to estimate region–label or event counts for operational envelopes.
$n_{\mathrm{cal}}$ Calibration sample size Number of points in $\mathcal{D}_{\mathrm{cal}}$.
$m$ Future (deployment) window size Number of future cases over which realized rates (and predictive envelopes) are defined.
Split conformal thresholding and SSBC
$s(x,y)$ Nonconformity score function Real-valued score for candidate label $y$ at input $x$; larger means “less conforming.”
$S_{i}:=s(X_{i},Y_{i})$ True-label calibration scores Scores computed on $\mathcal{D}_{\mathrm{cal}}$; exchangeable with the future score $S_{n+1}$.
$S_{(k)}$ $k$th order statistic of $\{S_{i}\}_{i=1}^{n_{\mathrm{cal}}}$ Used to define the deployed split-conformal threshold $\tau$.
$k$ Order-statistic index Split conformal threshold index; $k\in\{1,\dots,n_{\mathrm{cal}}\}$.
$u:=n_{\mathrm{cal}}+1-k$ Miscoverage grid index Equivalent parametrization of the conformal grid; $u\in\{1,\dots,n_{\mathrm{cal}}\}$.
$\alpha_{\mathrm{grid}}:=\dfrac{u}{n_{\mathrm{cal}}+1}$ Grid-aligned miscoverage level Miscoverage induced by selecting index $u$ on the discrete conformal grid.
$\tau:=S_{(k)}$ Deployed conformal threshold Fixed after calibration; determines the prediction-set map $C(\cdot)$.
$\alpha^{\star}$ Target (requested) miscoverage User-facing miscoverage level prior to discretization / SSBC selection.
$\delta$ Confidence (risk) parameter Tail probability for calibration-conditional under-coverage events (PAC-style semantics).
$p_{\mathrm{cov}}(\mathcal{D}_{\mathrm{cal}})$ Calibration-conditional coverage probability $p_{\mathrm{cov}}:=\mathbb{P}(Y\in C(X)\mid\mathcal{D}_{\mathrm{cal}})$; random over calibration draws. Under exchangeability, $p_{\mathrm{cov}}\sim\mathrm{Beta}(k,u)$ for split conformal.
$\widehat{C}_{m}$ Finite-window empirical coverage $\widehat{C}_{m}:=\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\{Y^{\prime}_{j}\in C(X^{\prime}_{j})\}$ over a future window of size $m$; $m=\infty$ denotes the infinite-window (calibration-conditional) regime.
Operational indicators, rates, and predictive envelopes
$C(X)$ Conformal prediction set Set-valued output of the deployed rule at input $X$.
$g_{\ell}(\cdot)$ Operational event functional Indicator map $g_{\ell}(C(X),Y)\in\{0,1\}$ defining KPI/event $\ell$ (e.g., singleton, wrong-singleton, hedged-positive).
$I_{\ell,i}$ Operational indicator (deployed rule) $I_{\ell,i}:=g_{\ell}(C(X_{i}),Y_{i})$ on an evaluation or deployment sample.
$K_{\ell}^{m}$ Future event count over a window $K_{\ell}^{m}:=\sum_{j=1}^{m}I_{\ell,j}$; the object governed by Binomial/Beta–Binomial predictive laws.
$\hat{R}_{\ell}:=\dfrac{1}{m}\sum_{j=1}^{m}I_{\ell,j}$ Realized operational rate over a window Empirical rate $\hat{R}_{\ell}=K_{\ell}^{m}/m$ over a window of size $m$.
Single-sample LOO surrogate (one dataset reuse)
$Z_{\ell,i}$ LOO operational indicator $Z_{\ell,i}:=g_{\ell}(C^{(-i)}(X_{i}),Y_{i})$ where $C^{(-i)}$ is calibrated on $\mathcal{D}_{\mathrm{cal}}\setminus\{(X_{i},Y_{i})\}$.
$N_{\mathrm{loo},\ell}:=\sum_{i=1}^{n_{\mathrm{cal}}}Z_{\ell,i}$ Pooled LOO indicator count LOO “success” count used to form LOO-based estimates and envelopes for event $\ell$; distinct from conformal $(k,u)$.
$\hat{r}_{\mathrm{loo},\ell}:=\dfrac{N_{\mathrm{loo},\ell}}{n_{\mathrm{cal}}}$ Pooled LOO rate estimate Convenience notation for the pooled LOO event frequency.
infl Inflation factor (controlled pessimization) Variance/width inflation used to hedge against dependence from single-sample reuse; implemented via $m_{\mathrm{eff}}=m/\texttt{infl}$.
Appendix B Exact finite-sample law of calibration-conditional coverage
This appendix derives an exact finite-sample characterization of the calibration-conditional coverage of a fixed split conformal predictor under exchangeability. The key point is that coverage is a pure rank event: the future true-label score is compared to a calibration order statistic. This yields a distribution-free pivot and an exact Beta law for realized (calibration-conditional) coverage across calibration draws. This pivot is the input to SSBC (Appendix C).
B.1 Setup
Let $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ be exchangeable with a future test pair $(X_{n+1},Y_{n+1})$. Let $s(x,y)$ be a nonconformity score and define
$S_{i}:=s(X_{i},Y_{i}),\quad i=1,\dots,n,\qquad\text{and}\qquad S_{n+1}:=s(X_{n+1},Y_{n+1}).$
Fix an index $k\in\{1,\dots,n\}$ and set
$\tau:=S_{(k)},\qquad u:=n+1-k,\qquad\alpha_{\mathrm{grid}}=\frac{u}{n+1}.$
For the predictor calibrated on $\mathcal{D}_{\mathrm{cal}}$, the calibration-conditional coverage probability is
$p_{\mathrm{cov}}(\mathcal{D}_{\mathrm{cal}}):=\mathbb{P}\!\left(S_{n+1}\leq\tau\mid\mathcal{D}_{\mathrm{cal}}\right).$
This random variable varies across calibration draws but is fixed once calibration is completed.
B.2 Rank pivot
Assume for clarity that the score distribution is continuous so ties occur with probability zero (ties are addressed in Remark 1). Under exchangeability of $\{S_{1},\dots,S_{n},S_{n+1}\}$, the rank
$R:=\operatorname{rank}\big(S_{n+1}\ \text{among}\ S_{1},\dots,S_{n},S_{n+1}\big)$
is uniform on $\{1,\dots,n+1\}$ (David and Nagaraja, 2003). Since $\tau=S_{(k)}$,
$\{S_{n+1}\leq\tau\}\quad\Longleftrightarrow\quad\{R\leq k\}.$
Thus coverage is the conditional probability of a rank event.
B.3 Exact distribution: Beta law
A standard order-statistic identity implies that the conditional probability of the rank event equals the $k$th order statistic of $n+1$ i.i.d. uniforms:
$p_{\mathrm{cov}}(\mathcal{D}_{\mathrm{cal}})\;\stackrel{{\scriptstyle d}}{{=}}\;U_{(k)},\qquad U_{(k)}\sim\mathrm{Beta}(k,u),\quad u=n+1-k,$
where $U_{(k)}$ is the $k$th order statistic of $U_{1},\dots,U_{n+1}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{Unif}(0,1)$ (David and Nagaraja, 2003). Equivalently, for any $t\in[0,1]$,
$\mathbb{P}\!\left(p_{\mathrm{cov}}(\mathcal{D}_{\mathrm{cal}})\leq t\right)=I_{t}(k,u),$
where $I_{t}(\cdot,\cdot)$ is the regularized incomplete Beta function. The law is distribution-free and depends only on $(n,k)$ (equivalently $(n,u)$).
Remark 1 (Discrete scores and ties).
If scores have atoms, ranks are not almost surely unique. Exact pivots can be recovered by randomized tie-breaking; deterministic left/right-continuous conventions yield conservative bounds. The non-interpolated order-statistic thresholding convention used in the main text preserves finite-sample validity.
Coverage depends only on the rank of the future true-label score relative to the calibration scores, hence admits an exact finite-sample distribution-free law. Most other operational KPIs depend on the joint geometry of multiple label scores and threshold interactions and therefore do not admit an analogous pivot (Appendix D).
Appendix C Small-Sample Beta Correction (SSBC)
SSBC is a deterministic index-selection rule for split conformal calibration that makes a user request $(\alpha^{\star},\delta)$ operationally precise for the single deployed predictor obtained after one calibration draw. SSBC selects a conformal grid index (equivalently an order statistic) so that, with probability at least $1-\delta$ over calibration randomness, the realized coverage of the deployed rule is at least $1-\alpha^{\star}$. In the finite-window variant, the same guarantee is imposed on empirical coverage over a future window of size $m$. The construction relies only on the exact finite-sample law of calibration-conditional coverage from Appendix B.
C.1 Setup and conformal grid
Let $n_{\mathrm{cal}}$ be the calibration size and let $S_{i}:=s(X_{i},Y_{i})$ denote true-label nonconformity scores for $i=1,\dots,n_{\mathrm{cal}}$. Split conformal selects a threshold as an order statistic:
$\tau=S_{(k)},\qquad k\in\{1,\dots,n_{\mathrm{cal}}\}.$
It is convenient to re-index the same grid by the miscoverage index
$u:=n_{\mathrm{cal}}+1-k\in\{1,\dots,n_{\mathrm{cal}}\},\qquad\alpha_{\mathrm{grid}}=\frac{u}{n_{\mathrm{cal}}+1},\qquad k=n_{\mathrm{cal}}+1-u.$
Here $\alpha^{\star}\in(0,1)$ is the requested miscoverage and $\delta\in(0,1)$ is a confidence/risk level controlling the probability (over calibration draws) that the requested semantics fail.
C.2 Exact law of realized coverage
For the fixed predictor calibrated on $\mathcal{D}_{\mathrm{cal}}$, define the calibration-conditional coverage probability
$p_{\mathrm{cov}}(\mathcal{D}_{\mathrm{cal}}):=\mathbb{P}(S_{n+1}\leq\tau\mid\mathcal{D}_{\mathrm{cal}}),$
where $(X_{n+1},Y_{n+1})$ is exchangeable with the calibration sample. Under exchangeability, Appendix B shows that
$p_{\mathrm{cov}}(\mathcal{D}_{\mathrm{cal}})\sim\mathrm{Beta}(k,u),\qquad k=n_{\mathrm{cal}}+1-u,$
exactly and distribution-free. This law describes how the realized coverage of the deployed predictor varies across hypothetical recalibrations, while treating the deployed rule as fixed after the one calibration step.
C.3 SSBC objective: calibration-conditional PAC semantics
SSBC enforces the calibration-conditional PAC-style constraint
$\mathbb{P}_{\mathcal{D}_{\mathrm{cal}}}\!\left(p_{\mathrm{cov}}(\mathcal{D}_{\mathrm{cal}})\ \geq\ 1-\alpha^{\star}\right)\ \geq\ 1-\delta.$
Using the Beta law, this is equivalent to
$\mathbb{P}\!\left(Z\geq 1-\alpha^{\star}\right)\ \geq\ 1-\delta,\qquad Z\sim\mathrm{Beta}(k,u).$
Selection rule: least conservative admissible grid point.
Among discrete grid indices $u\in\{1,\dots,n_{\mathrm{cal}}\}$, SSBC selects the largest admissible $u$ satisfying the tail constraint above:
$u^{\star}:=\max\Big\{u\in\{1,\dots,n_{\mathrm{cal}}\}:\ \mathbb{P}(Z\geq 1-\alpha^{\star})\geq 1-\delta,\ Z\sim\mathrm{Beta}(n_{\mathrm{cal}}+1-u,\ u)\Big\}.$
Equivalently, SSBC deploys the least conservative grid miscoverage level $\alpha_{\mathrm{adj}}=u^{\star}/(n_{\mathrm{cal}}+1)$ that certifies the requested semantics. The returned order-statistic index is $k_{\mathrm{adj}}=n_{\mathrm{cal}}+1-u^{\star}$.
C.4 Finite-window deployment semantics
In many deployments, coverage is evaluated over a finite window of size $m$. Define empirical coverage over that window by
$\widehat{C}_{m}:=\frac{1}{m}\sum_{j=1}^{m}I_{\mathrm{cov},j},\qquad I_{\mathrm{cov},j}:=\mathbf{1}\{Y^{\prime}_{j}\in C(X^{\prime}_{j})\},\qquad S_{m}:=m\widehat{C}_{m}.$
Conditional on the calibration-conditional coverage probability $p_{\mathrm{cov}}(\mathcal{D}_{\mathrm{cal}})=p$,
$S_{m}\mid p\sim\mathrm{Binomial}(m,p).$
Marginalizing $p\sim\mathrm{Beta}(k,u)$ yields the exact mixture law
$S_{m}\sim\mathrm{Beta\text{-}Binomial}(m;k,u),\qquad k=n_{\mathrm{cal}}+1-u.$
The finite-window SSBC criterion selects $u$ such that
$\mathbb{P}\!\left(\widehat{C}_{m}\geq 1-\alpha^{\star}\right)\ \geq\ 1-\delta,$
where the probability is over both calibration randomness and the future window.
C.4.1 Strict lower-tail convention
Because $\widehat{C}_{m}$ is discrete, we adopt a strict violation convention $\widehat{C}_{m}<1-\alpha^{\star}$. Define the corresponding count threshold
$x^{\star}:=\left\lfloor(1-\alpha^{\star})m\right\rfloor+1,\qquad\text{so that}\qquad\{\widehat{C}_{m}\geq 1-\alpha^{\star}\}\iff\{S_{m}\geq x^{\star}\}.$
All Beta–Binomial tail probabilities in SSBC are evaluated for the event $\{S_{m}\geq x^{\star}\}$. This avoids boundary ambiguity when $(1-\alpha^{\star})m$ is an integer and preserves monotonicity in $u$.
C.5 Infinite-window limit
This limiting regime is identified by Marques (2025). In our notation, if $u_{m}$ denotes the SSBC-selected grid index obtained by inverting the Beta–Binomial tail for window size $m$, and $u_{\infty}$ denotes the index obtained from the Beta limit law, then $u_{m}\to u_{\infty}$ as $m\to\infty$. Intuitively, conditional on $p_{\mathrm{cov}}=p$, the empirical coverage $\widehat{C}_{m}$ converges almost surely to $p$, so the finite-window Beta–Binomial criterion approaches the infinite-window Beta criterion.
C.6 Feasibility and saturation
Not every pair $(\alpha^{\star},\delta)$ is feasible at fixed $n_{\mathrm{cal}}$ because calibration choices lie on the conformal grid. Under the most conservative grid point $u=1$ (equivalently $k=n_{\mathrm{cal}}$),
$p_{\mathrm{cov}}\sim\mathrm{Beta}(n_{\mathrm{cal}},1),\qquad\mathbb{P}\!\left(p_{\mathrm{cov}}\geq 1-\alpha\right)=1-(1-\alpha)^{n_{\mathrm{cal}}}.$
Thus any infinite-window PAC requirement at confidence $1-\delta$ must satisfy
$\alpha\ \geq\ 1-\delta^{1/n_{\mathrm{cal}}}.$
If this condition is violated (and likewise in the finite-window analogue), no grid point can satisfy the tail constraint and SSBC returns Infeasible.
C.7 Coverage semantics under finite calibration
To visualize how nominal requests map to deployed semantics under finite calibration, Figure 6 plots the effective calibration level $\alpha_{\mathrm{adj}}$ selected by SSBC as a function of the user inputs $(\alpha,\delta)$ at fixed $n_{\mathrm{cal}}$. Distinct nominal requests can induce the same deployed grid index and therefore the same realized coverage semantics.
Figure 6: Semantic interpretation of nominal coverage requests under finite calibration.Each panel visualizes the effective calibration level $\alpha_{\mathrm{adj}}$ selected by SSBC as a function of the user-specified miscoverage level $\alpha$ and confidence level $\delta$, for fixed $n_{\mathrm{cal}}$. Color encodes the deployed level $\alpha_{\mathrm{adj}}$, while contours indicate iso-semantic sets: distinct nominal requests that induce the same deployed calibration grid point. The feasibility boundary reflects the finite-sample constraint $\alpha\gtrsim 1-\delta^{1/n_{\mathrm{cal}}}$.
C.8 SSBC algorithm (deterministic specification)
This subsection provides a reproducible implementation-level specification. The algorithm returns the largest admissible $u$ (least conservative grid point) satisfying the relevant tail constraint.
Algorithm 1 Small-Sample Beta Correction (SSBC)
1:Target miscoverage $\alpha^{\star}\in(0,1)$; calibration size $n_{\mathrm{cal}}\in\mathbb{N}$; confidence $\delta\in(0,1)$; deployment regime $\in\{\infty,m\}$ (window size $m$ if finite)
2:Adjusted grid level $\alpha_{\mathrm{adj}}$ and index $k_{\mathrm{adj}}$, or Infeasible
[…]
5:if regime =m=m then
6:  $x^{\star}\leftarrow\lfloor t\,m\rfloor+1$
7:end if
8:for $u=1,\dots,n_{\mathrm{cal}}$ do
9:  $a\leftarrow n_{\mathrm{cal}}+1-u$,  $b\leftarrow u$
10:  if regime $\infty$ then
11:   $p_{\mathrm{tail}}\leftarrow\Pr[Z\geq t],\ Z\sim\mathrm{Beta}(a,b)$
12:  else
13:   $p_{\mathrm{tail}}\leftarrow\Pr[X\geq x^{\star}],\ X\sim\mathrm{Beta\text{-}Binomial}(m;a,b)$
14:  end if
15:  if $p_{\mathrm{tail}}\geq 1-\delta$ then
16:   $u^{\star}\leftarrow u$
17:  end if
18:end for
19:if $u^{\star}<0$ then
20:  return Infeasible
21:end if
[…]
24:return $\alpha_{\mathrm{adj}},\ k_{\mathrm{adj}}$
Relation to DKWM-style calibration.
DKWM-style calibration modifies the nominal grid choice to enforce conservative, worst-case guarantees uniformly over calibration draws and distributions. SSBC addresses a different question: it assigns a calibration-conditional PAC meaning to a user request $(\alpha^{\star},\delta)$ for the single deployed predictor produced after one calibration. DKWM targets uniform validity across hypothetical recalibrations; SSBC targets admissibility of the realized rule via exact Beta (or Beta–Binomial) tails.
C.9 Extended simulation table for calibration-conditional violations
Table 7 reports the full calibration-size grid for the simulation summarized in Section 5.1.1.
Table 7: Calibration-conditional coverage violation rates with theory. Target miscoverage $\alpha^{\star}=0.10$, confidence $\delta=0.10$. $\alpha_{\mathrm{grid}}$ is the grid point selected on the conformal grid, $\alpha_{\mathrm{cont}}$ is the requested value under the DKWM correction, and $m_{\mathrm{cal}}$ is the calibration-window size. Results are based on $10^{6}$ calibration draws and a finite deployment window of size $m_{\mathrm{infer}}=100$. The Obs column reports $\mathbb{P}(\widehat{C}_{m}<1-\alpha^{\star})$. Beta reports $\mathbb{P}(p_{\mathrm{cov}}<1-\alpha^{\star})$ under $p_{\mathrm{cov}}\sim\mathrm{Beta}(k,u)$. BetaBinom reports $\mathbb{P}(\widehat{C}_{m}<1-\alpha^{\star})$ under the induced Beta–Binomial law.
$n_{\mathrm{cal}}$ Method $u$ $\alpha_{\mathrm{grid}}$ Obs Beta BetaBinom $\alpha_{\mathrm{cont}}$
50 None 5 0.0980 0.3963 0.4312 0.3964
50 SSBC 2 0.0392 0.0476 0.0338 0.0472
50 DKWM 1 0.0196 0.0096 0.0052 0.0095 0.0731-0.0731
75 None 7 0.0921 0.3454 0.3673 0.3464
75 SSBC 4 0.0526 0.0769 0.0504 0.0768
75 DKWM 1 0.0132 0.0016 0.0004 0.0017 0.0413-0.0413
100 None 10 0.0990 0.4075 0.4513 0.4071
100 SSBC 6 0.0594 0.0960 0.0576 0.0956
100 DKWM 1 0.0099 0.0004 0.0000 0.0004 0.0224-0.0224
150 None 15 0.0993 0.4119 0.4602 0.4107
150 SSBC 9 0.0596 0.0804 0.0307 0.0801
150 DKWM 1 0.0066 0.0000 0.0000 0.0000 +0.0001+0.0001
200 None 20 0.0995 0.4130 0.4655 0.4124
200 SSBC 13 0.0647 0.0974 0.0320 0.0980
200 DKWM 2 0.0100 0.0000 0.0000 0.0000 +0.0135+0.0135
250 None 25 0.0996 0.4126 0.4692 0.4134
250 SSBC 16 0.0637 0.0863 0.0175 0.0858
250 DKWM 5 0.0199 0.0003 0.0000 0.0003 +0.0226+0.0226
300 None 30 0.0997 0.4139 0.4719 0.4141
300 SSBC 20 0.0664 0.0969 0.0171 0.0971
300 DKWM 8 0.0266 0.0009 0.0000 0.0009 +0.0293+0.0293
500 None 50 0.0998 0.4153 0.4782 0.4153
500 SSBC 34 0.0679 0.0949 0.0049 0.0955
500 DKWM 22 0.0439 0.0097 0.0000 0.0096 +0.0453+0.0453
Appendix D Single-sample structural coupling, leave-one-out decoupling, and envelope inflation
The two-stage predictive reference used throughout this paper separates calibration (choose a conformal threshold on $\mathcal{D}_{\mathrm{cal}}$) from operational evaluation (estimate rates on an independent window). This separation is what makes window indicators behave as i.i.d. Bernoulli draws under a fixed deployed rule.
In practice, an independent audit set is not always be available. When the same sample is reused both to select the threshold and to estimate operational rates, threshold selection and evaluation become coupled. This appendix records (i) a minimal structural reason for reuse-induced dependence, (ii) a data-efficient remedy—leave-one-out (LOO) recalibration—that provides effective practical decoupling in our regimes, and (iii) an inflation parameter infl that pessimizes predictive envelopes when residual dependence or regime instability remains.
D.1 Why single-sample reuse introduces dependence
Let $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ be an exchangeable calibration sample, let $S_{i}:=s(X_{i},Y_{i})$ be true-label nonconformity scores, and let the split conformal threshold be the $k$th order statistic
$\hat{\tau}:=S_{(k)},\qquad k\in\{1,\dots,n\},$
assuming continuity so ties occur with probability zero. For any threshold $t$, define the crossing indicator $I_{i}(t):=\mathbf{1}\{S_{i}\leq t\}$.
Insight from counting.
If $\hat{\tau}=t$ is the $k$th order statistic, then exactly k1k-1 scores lie strictly below $t$ and one score equals $t$ (under continuity, almost surely), so
$\sum_{i=1}^{n}I_{i}(t)=k\quad\text{almost surely}.$
This already shows that the indicators $\{I_{i}(t)\}_{i=1}^{n}$ cannot be conditionally independent given $\hat{\tau}=t$: once some are known to be one, fewer ones remain available for the others. By exchangeability, for any $i\neq j$,
$\mathrm{Cov}\!\left(I_{i}(t),I_{j}(t)\mid\hat{\tau}=t\right):=\frac{k(k-n)}{n^{2}(n-1)}<0\quad(k<n).$
Implication for operational envelopes.
Operational indicators are functions of the deployed prediction set $C_{\hat{\tau}}(X)=\{y:s(X,y)\leq\hat{\tau}\}$ and therefore inherit reuse-induced dependence. In particular, the nonzero conditional covariance shows that these indicators are not Bernoulli draws under a fixed rule once threshold selection and evaluation are performed on the same sample. To recover the Bernoulli sampling picture needed for the two-stage guarantees in the main text, one therefore needs an independent audit set after calibration. If one naïvely treats reuse-based indicators as i.i.d. Bernoulli under a fixed rule, predictive envelopes can become under-dispersed. Empirically in our studies, the dominant distortion is in dispersion: point estimates often remain close to a two-stage reference, while intervals can be too narrow without decoupling or pessimization.
D.2 Approximate decoupling via leave-one-out (LOO)
When an independent audit sample is unavailable, we use LOO recalibration (Vovk, 2015; Barber et al., 2021) to reduce self-influence. Let $\hat{\tau}_{-i}$ be the split conformal threshold computed on $\mathcal{D}_{\mathrm{cal}}\setminus\{(X_{i},Y_{i})\}$, and let $C_{-i}(\cdot)$ be the corresponding prediction set map. For an operational event functional $g_{j}(C(X),Y)\in\{0,1\}$ (e.g., singleton, doublet, wrong-singleton), define LOO indicators
$Z_{i,j}:=g_{j}\!\left(C_{-i}(X_{i}),Y_{i}\right),\qquad i=1,\dots,n,$
and pooled LOO summaries
$k_{\mathrm{pool},j}:=\sum_{i=1}^{n}Z_{i,j},\qquad\widehat{r}^{\mathrm{LOO}}_{j}:=\frac{k_{\mathrm{pool},j}}{n}.$
Proxy two-stage interpretation.
Each $Z_{i,j}$ is evaluated under a rule that does not use point $i$, restoring a localized separation between rule construction and evaluation. Although the rule varies across folds, each fold differs only slightly from the full calibrated rule, so $\{Z_{i,j}\}$ can be read as indicators from nearby operating regimes. Pooling provides a direct empirical proxy for operational behavior under finite calibration.
Empirical decoupling.
In the regimes studied (Section 5.1.2), LOO envelopes track the two-stage Calibrate–and–Audit reference closely in center and often slightly pessimistically in width. We therefore use pooled LOO indicators for planning when a separate audit set is unavailable.
D.3 Envelope inflation as controlled pessimization
LOO is local and does not remove all dependence, especially near regime boundaries where small threshold shifts can change region support. We therefore introduce an explicit inflation parameter $\texttt{infl}\geq 1$ that widens predictive envelopes by shrinking an effective sample size used in the predictive model.
Operational effect.
In constructions below we replace a nominal proxy sample size $n$ by $n_{\mathrm{eff}}=n/\texttt{infl}$ (and analogously for any proxy count used to parameterize predictive dispersion). Larger infl yields wider envelopes without changing the pooled mean, providing a monotone knob for conservatism.
Diagnostic guidance.
As a simple diagnostic of calibration-induced variability, one can compute fold rates $\widehat{r}^{(-i)}$ under each LOO-calibrated rule and their empirical variance
$\overline{r}_{\mathrm{LOO}}=\frac{1}{n}\sum_{i=1}^{n}\widehat{r}^{(-i)},\qquad\widehat{\mathrm{Var}}_{\mathrm{LOO}}=\frac{1}{n-1}\sum_{i=1}^{n}\left(\widehat{r}^{(-i)}-\overline{r}_{\mathrm{LOO}}\right)^{2}.$
In our experiments this variance aligns qualitatively with inflation levels needed for conservative envelope coverage relative to the two-stage reference.
D.4 Predictive envelope constructions from LOO proxies
We describe two complementary envelope constructions built from LOO proxy indicators. The first mirrors the two-stage reference and is our primary approximation; the second is a conservative guardrail.
Predictive Beta–Binomial envelopes.
Let $Z_{i}$ denote a pooled proxy indicator for a fixed KPI (suppress $j$), and let $k_{\mathrm{pool}}=\sum_{i=1}^{n}Z_{i}$ with pooled proxy rate $\widehat{p}=k_{\mathrm{pool}}/n$. Apply inflation via
$n_{\mathrm{eff}}=\frac{n}{\texttt{infl}},\qquad k_{\mathrm{eff}}=\widehat{p}\,n_{\mathrm{eff}}=\frac{k_{\mathrm{pool}}}{\texttt{infl}}.$
With a small prior offset $\mathrm{offset}\in\{1,1/2\}$, define
$\alpha:=k_{\mathrm{eff}}+\mathrm{offset},\qquad\beta:=(n_{\mathrm{eff}}-k_{\mathrm{eff}})+\mathrm{offset}.$
For a future deployment window of size $m$, the predictive count $S_{m}$ is modeled as
$S_{m}\sim\mathrm{BetaBinomial}(m,\alpha,\beta),$
and equal-tailed prediction intervals follow from the Beta–Binomial CDF. Increasing infl shrinks $n_{\mathrm{eff}}$ and widens intervals monotonically while preserving the pooled mean.
Hoeffding-type dominance bound.
As a conservative alternative, Hoeffding’s inequality gives a distribution-free bound for a window rate $\widehat{r}_{m}$ around a proxy mean $\widehat{r}_{\mathrm{LOO}}$:
$\Pr\!\left(\left|\widehat{r}_{m}-\widehat{r}_{\mathrm{LOO}}\right|\geq\epsilon\right)\leq 2\exp(-2m\epsilon^{2}),$
yielding simple symmetric envelopes that can be used as a worst-case planning check.
Summary.
Single-sample reuse induces structural dependence through the thresholding mechanism described above, primarily distorting predictive dispersion. LOO recalibration reduces self-influence and empirically recovers audit-style behavior in our regimes. Predictive envelopes are then constructed from pooled LOO proxies using a Beta–Binomial model, with infl providing a monotone knob for controlled pessimization and the Hoeffding bound serving as a conservative guardrail.
Appendix E Binary conformal partitions: regimes, coupling, and rate primitives
This appendix records the geometric facts used in Section 2, Section 3, and Section 4. Here we assume thresholds are fixed and all probabilities are taken with respect to the deployment distribution conditional on $\mathcal{D}_{\mathrm{cal}}$.
E.1 Class-conditional split conformal as a four-region partition.
Let $\mathcal{Y}=\{0,1\}$ and let $s(x,y)$ be a nonconformity score. Class-conditional split conformal produces thresholds $\tau_{0},\tau_{1}$, and the set-valued output is
$\mathcal{C}(x)=\{0:s(x,0)\leq\tau_{0}\}\cup\{1:s(x,1)\leq\tau_{1}\},$ (2)
equivalently represented by the region label
$R_{\tau}(x):=\big(\mathbf{1}\{s(x,0)\leq\tau_{0}\},\ \mathbf{1}\{s(x,1)\leq\tau_{1}\}\big)\in\{0,1\}^{2},\qquad\tau=(\tau_{0},\tau_{1}).$
Writing $(s_{0},s_{1})=(s(x,0),s(x,1))$, the thresholds partition score space into
$R_{\tau}(x)=\begin{cases}11,&s_{0}\leq\tau_{0},\ s_{1}\leq\tau_{1}\quad\text{(doublet)},\\ 10,&s_{0}\leq\tau_{0},\ s_{1}>\tau_{1}\quad\text{(singleton $\{0\}$)},\\ 01,&s_{0}>\tau_{0},\ s_{1}\leq\tau_{1}\quad\text{(singleton $\{1\}$)},\\ 00,&s_{0}>\tau_{0},\ s_{1}>\tau_{1}\quad\text{(abstention)}.\end{cases}$
For any fixed $\tau$,
$\sum_{r\in\{00,01,10,11\}}\mu_{r}(\tau)=1,\qquad\mu_{r}(\tau):=\Pr(R_{\tau}(X)=r\mid\mathcal{D}_{\mathrm{cal}}).$
E.2 Probability-normalized scores and a sharp regime boundary
Many probabilistic classifiers induce probability-normalized scores, e.g. $s(x,y)=1-P(y\mid x)$. For $\mathcal{Y}=\{0,1\}$ this implies
$s(x,0)+s(x,1)=1,$
so feasible score pairs lie on the diagonal manifold
$\mathcal{M}=\{(u,1-u):u\in[0,1]\}.$
Intersecting $\mathcal{M}$ with the threshold rectangles yields a sharp boundary that determines which region types can occur with nonzero mass.
Proposition 2 (Regime boundary under probability normalization).
Assume $(s_{0},s_{1})\in\mathcal{M}$ almost surely. Then:
  1. 1.
    $R_{11}$ has nonempty intersection with $\mathcal{M}$ iff $\tau_{0}+\tau_{1}\geq 1$, and has positive-length intersection iff $\tau_{0}+\tau_{1}>1$.
  2. 2.
    $R_{00}$ has nonempty intersection with $\mathcal{M}$ iff $\tau_{0}+\tau_{1}\leq 1$, and has positive-length intersection iff $\tau_{0}+\tau_{1}<1$.
  3. 3.
    On the boundary $\tau_{0}+\tau_{1}=1$, both $R_{11}$ and $R_{00}$ intersect $\mathcal{M}$ at a single point; hence under any continuous distribution on $\mathcal{M}$ they have probability zero and only the singleton regions carry mass.
Proof.
Parameterize $\mathcal{M}$ by $u=s_{0}\in[0,1]$, so $s_{1}=1-u$.
$R_{11}$ requires $u\leq\tau_{0}$ and $1-u\leq\tau_{1}$, i.e. $u\in[\,1-\tau_{1},\ \tau_{0}\,]$. This interval has positive length iff $1-\tau_{1}<\tau_{0}$, i.e. $\tau_{0}+\tau_{1}>1$, and degenerates to a point at equality.
$R_{00}$ requires $u>\tau_{0}$ and $1-u>\tau_{1}$, i.e. $u\in(\,\tau_{0},\ 1-\tau_{1}\,)$. This interval has positive length iff $\tau_{0}<1-\tau_{1}$, i.e. $\tau_{0}+\tau_{1}<1$, and degenerates to a point at equality. ∎
Crossing the affine boundary $\tau_{0}+\tau_{1}=1$ therefore removes an entire region label (doublet or abstention) from the support of $R_{\tau}(X)$ under probability normalization, explaining sharp regime changes in attainable operating behavior.
E.3 Cross-threshold dominance in the hedging regime
Within the hedging regime $\tau_{0}+\tau_{1}>1$, the manifold $\mathcal{M}$ is partitioned into three contiguous intervals corresponding to $\{10,11,01\}$. With $u=s(x,0)$:
$u\in[0,\,1-\tau_{1})\Rightarrow R_{\tau}(x)=10,\qquad u\in[\,1-\tau_{1},\,\tau_{0}\,]\Rightarrow R_{\tau}(x)=11,\qquad u\in(\tau_{0},\,1]\Rightarrow R_{\tau}(x)=01.$
Hence each singleton region is controlled primarily by the opposing threshold: increasing $\tau_{1}$ expands $11$ at the expense of $10$, while increasing $\tau_{0}$ expands $11$ at the expense of $01$. Operationally, $(\tau_{0},\tau_{1})$ act as mass-reallocation boundaries, not independent per-class knobs.
E.4 Region–label primitives and rate factorizations
For auditing and planning, the primitive object is the region–label table
$p_{r,y}(\tau):=\Pr(R_{\tau}(X)=r,\ Y=y\mid\mathcal{D}_{\mathrm{cal}}),\qquad r\in\{00,01,10,11\},\ y\in\{0,1\}.$
Two derived summaries are
$\mu_{r}(\tau)$ $:=\Pr(R_{\tau}(X)=r\mid\mathcal{D}_{\mathrm{cal}})=\sum_{y\in\{0,1\}}p_{r,y}(\tau),$
$\eta_{r}(\tau)$ $:=\Pr(Y=1\mid R_{\tau}(X)=r,\ \mathcal{D}_{\mathrm{cal}})=\frac{p_{r,1}(\tau)}{\mu_{r}(\tau)}\quad(\mu_{r}(\tau)>0).$
Any region-measurable KPI is a projection of $\{p_{r,y}(\tau)\}$; see Section 2.4.
Example: decisive error masses under “commit-on-singletons”.
Consider the convention
$10\mapsto 0,\qquad 01\mapsto 1,\qquad 11,00\mapsto\text{defer}.$
Then the decisive false-negative and false-positive masses are
$\mathrm{FN}_{\mathrm{dec}}(\tau)=\Pr(R_{\tau}(X)=10,\ Y=1\mid\mathcal{D}_{\mathrm{cal}})=p_{10,1}(\tau)=\mu_{10}(\tau)\,\eta_{10}(\tau),$
$\mathrm{FP}_{\mathrm{dec}}(\tau)=\Pr(R_{\tau}(X)=01,\ Y=0\mid\mathcal{D}_{\mathrm{cal}})=p_{01,0}(\tau)=\mu_{01}(\tau)\,[1-\eta_{01}(\tau)].$
The decisive mass is $\mu_{10}(\tau)+\mu_{01}(\tau)$, while the defer mass is $\mu_{11}(\tau)+\mu_{00}(\tau)$ (with feasibility of $11$ versus $00$ governed by Proposition 2 under probability normalization).
Appendix F Cost-coherence and inverse pricing envelopes induced by a fixed conformal interface
This appendix formalizes the operational point that distribution-free calibration is not automatically cost-agnostic once conformal outputs are wired into actions. Fix thresholds $\tau$ (calibration-conditional viewpoint) and treat the induced finite region label $R_{\tau}(X)\in\mathcal{R}$ as the deployed observable. Any downstream rule that uses only this interface can depend on the data only through $r\in\mathcal{R}$, so whether the convention is coherent with a cost model is evaluated region-wise using within-region label frequencies (we formalize this as cost-coherence below). We cast this as an inverse pricing problem: given a fixed (interface, convention) pair, characterize the set of consequence prices under which the convention is coherent.
F.1 Interface primitives: masses and within-region label frequencies
Specialize to $\mathcal{Y}=\{0,1\}$ and $\mathcal{R}=\{00,01,10,11\}$ as in Appendix E. Adopt the calibration-conditional joint table
$p_{r,y}:=\Pr(R_{\tau}(X)=r,\ Y=y\mid\mathcal{D}_{\mathrm{cal}}).$
Define region mass and within-region label frequency
$\mu_{r}:=p_{r,0}+p_{r,1},\qquad\eta_{r}:=\Pr(Y=1\mid R_{\tau}(X)=r,\ \mathcal{D}_{\mathrm{cal}})=\frac{p_{r,1}}{\mu_{r}}\quad(\mu_{r}>0).$
Because $\mathcal{R}$ is finite, conditional uncertainty about $Y$ is piecewise constant: within region $r$, all decision comparisons reduce to $\eta_{r}$.
F.2 Region-associated action conventions and coherence
Let $\mathcal{A}$ be a finite action set and let $L_{\theta}(a,y)$ be a priced consequence (cost or negative utility), parameterized by $\theta\in\Theta$. A deployed convention is a region-associated policy
$\tilde{\pi}:\mathcal{R}\to\mathcal{A}.$
For region $r$ with $\mu_{r}>0$, the interface-relative conditional risk of action $a$ is
$\mathcal{C}_{\theta}(a\mid r):=\mathbb{E}[L_{\theta}(a,Y)\mid R_{\tau}(X)=r,\ \mathcal{D}_{\mathrm{cal}}]=(1-\eta_{r})L_{\theta}(a,0)+\eta_{r}L_{\theta}(a,1).$
Cost-coherence (interface-relative optimality).
We say $\tilde{\pi}$ is cost-coherent on the deployed interface under pricing $\theta$ if for every region with $\mu_{r}>0$,
$\mathcal{C}_{\theta}(\tilde{\pi}(r)\mid r)\leq\mathcal{C}_{\theta}(a\mid r)\qquad\forall a\in\mathcal{A}.$
This is the dominance condition induced by acting only on $R_{\tau}(X)$.
F.3 Inverse pricing envelope
For each region with $\mu_{r}>0$, define the local feasibility set
$\Theta_{r}(\tilde{\pi}):=\Bigl\{\theta\in\Theta:\ \mathcal{C}_{\theta}(\tilde{\pi}(r)\mid r)\leq\mathcal{C}_{\theta}(a\mid r)\ \ \forall a\in\mathcal{A}\Bigr\},$
and the global pricing envelope
[…]
F.4 Worked case: Chow-style reject option and “commit on singletons”
Consider $\mathcal{A}=\{0,1,\mathrm{rej}\}$ with Chow-style costs (false negative $c_{01}>0$, false positive $c_{10}>0$, rejection $c_{\mathrm{rej}}\geq 0$):
$L(0,1)=c_{01},\quad L(1,0)=c_{10},\quad L(\mathrm{rej},y)=c_{\mathrm{rej}},\quad L(0,0)=L(1,1)=0.$
In region $r$ the conditional risks are
$\mathcal{C}(0\mid r)=\eta_{r}\,c_{01},\qquad\mathcal{C}(1\mid r)=(1-\eta_{r})\,c_{10},\qquad\mathcal{C}(\mathrm{rej}\mid r)=c_{\mathrm{rej}}.$
Work in ratios
$\lambda:=\frac{c_{01}}{c_{10}},\qquad\rho:=\frac{c_{\mathrm{rej}}}{c_{10}},$
so only $(\lambda,\rho)$ matters for comparisons.
Convention.
$\tilde{\pi}(10)=0,\qquad\tilde{\pi}(01)=1,\qquad\tilde{\pi}(11)=\mathrm{rej},\qquad\tilde{\pi}(00)=\mathrm{rej}.$
Coherence holds iff the region-wise dominance inequalities below hold for all regions with $\mu_{r}>0$.
Singleton region $01$ (output $\{1\}$).
Choosing $1$ must beat both 0 and $\mathrm{rej}$:
$(1-\eta_{01})\leq\eta_{01}\lambda\quad\Longleftrightarrow\quad\eta_{01}\geq\frac{1}{1+\lambda},\qquad(1-\eta_{01})\leq\rho\quad\Longleftrightarrow\quad\eta_{01}\geq 1-\rho.$
Thus
$\eta_{01}\ \geq\ \max\!\left\{\frac{1}{1+\lambda},\ 1-\rho\right\}.$
Singleton region $10$ (output $\{0\}$).
Choosing 0 must beat both $1$ and $\mathrm{rej}$:
$\eta_{10}\lambda\leq(1-\eta_{10})\quad\Longleftrightarrow\quad\eta_{10}\leq\frac{1}{1+\lambda},\qquad\eta_{10}\lambda\leq\rho\quad\Longleftrightarrow\quad\eta_{10}\leq\frac{\rho}{\lambda}.$
Thus
$\eta_{10}\ \leq\ \min\!\left\{\frac{1}{1+\lambda},\ \frac{\rho}{\lambda}\right\}.$
Rejection regions $r\in\{11,00\}$.
Rejecting must beat both commitments:
$\rho\leq\eta_{r}\lambda\quad\Longleftrightarrow\quad\eta_{r}\geq\frac{\rho}{\lambda},\qquad\rho\leq(1-\eta_{r})\quad\Longleftrightarrow\quad\eta_{r}\leq 1-\rho.$
Hence rejection is optimal on region $r$ only if
$\frac{\rho}{\lambda}\ \leq\ \eta_{r}\ \leq\ 1-\rho.$
The rejection band is nonempty only if
$\frac{\rho}{\lambda}\leq 1-\rho\quad\Longleftrightarrow\quad\rho\leq\frac{\lambda}{1+\lambda}\quad\Longleftrightarrow\quad c_{\mathrm{rej}}\leq\frac{c_{01}c_{10}}{c_{01}+c_{10}}.$
[…]
Appendix G Explicit region indicators and projection masks
This appendix briefly instantiates the linear “sum selected cells” formalism from Section 2.4 in the binary toy geometry used in Figure 2 and Figure 3 ($R_{\tau}(x)\xrightarrow{\ \pi\ }C(x)$). Region feasibility and regime facts are recorded in Appendix E; here we keep only the explicit projection masks needed for audit computations.
G.1 Region labels and the region–label table
Assume $\mathcal{Y}=\{0,1\}$ and thresholds $\tau=(\tau_{0},\tau_{1})$. Define the deployed region label
$R_{\tau}(x)=(r_{0}(x),r_{1}(x))\in\{0,1\}^{2},\qquad r_{y}(x):=\mathbf{1}\{s(x,y)\leq\tau_{y}\},$
with region names $\mathcal{R}=\{r_{10},r_{11},r_{01},r_{00}\}$ as in Appendix E.1. For a calibration setting $\theta$ (indexing deployed thresholds $\tau(\theta)$), define the calibration-conditional region–label table
$P(\theta)=\big(p_{r,y}(\theta)\big)_{r\in\mathcal{R},\ y\in\mathcal{Y}},\qquad p_{r,y}(\theta)=\Pr\!\big(R_{\tau(\theta)}(X)=r,\ Y=y\mid\mathcal{D}_{\mathrm{cal}}\big).$
A linear operational rate is obtained by summing the subset of cells $(r,y)$ that define the event.
G.2 Policies used in Figure 3
A policy $\pi$ maps region labels to prediction sets $C(x)\subseteq\{0,1\}$:
$R_{\tau}(x)\ \xrightarrow{\ \pi\ }\ C(x).$
The three policies used in Figure 3 are:
Set inclusion $\pi_{\mathrm{SI}}$.
$\pi_{\mathrm{SI}}(r_{10})=\{0\},\quad\pi_{\mathrm{SI}}(r_{11})=\{0,1\},\quad\pi_{\mathrm{SI}}(r_{01})=\{1\},\quad\pi_{\mathrm{SI}}(r_{00})=\varnothing.$
Commit–reject $\pi_{\mathrm{CR}}$.
$\pi_{\mathrm{CR}}(r_{10})=\{0\},\quad\pi_{\mathrm{CR}}(r_{01})=\{1\},\quad\pi_{\mathrm{CR}}(r_{11})=\varnothing,\quad\pi_{\mathrm{CR}}(r_{00})=\varnothing.$
Set exclusion $\pi_{\mathrm{SE}}$ (complement of set inclusion).
$\pi_{\mathrm{SE}}(r_{10})=\{1\},\quad\pi_{\mathrm{SE}}(r_{11})=\varnothing,\quad\pi_{\mathrm{SE}}(r_{01})=\{0\},\quad\pi_{\mathrm{SE}}(r_{00})=\{0,1\}.$
G.3 Binary projection masks
Fix a policy $\pi$. For any event/quantity $\ell$, define a $4\times 2$ binary mask
$G_{\ell}(\pi)=\big(g_{\ell}(r,y;\pi)\big)_{r\in\mathcal{R},\ y\in\mathcal{Y}}\in\{0,1\}^{4\times 2},$
where $g_{\ell}(r,y;\pi)=1$ indicates that the cell $(r,y)$ is included in the sum. Then the corresponding rate is
$r_{\ell}(\theta)=\sum_{r\in\mathcal{R}}\sum_{y\in\mathcal{Y}}g_{\ell}(r,y;\pi)\,p_{r,y}(\theta).$
We record one detailed worked example and then list several shorter special cases used elsewhere in the paper.
(1) Coverage under set inclusion.
Under $\pi_{\mathrm{SI}}$, coverage is the event $\{Y\in C(X)\}$. Cell-by-cell, coverage holds for: (i) $y=0$ in regions $r_{10}$ or $r_{11}$, and (ii) $y=1$ in regions $r_{01}$ or $r_{11}$. Hence
$G_{\mathrm{cov}}(\pi_{\mathrm{SI}})=\begin{array}[]{c|cc}&y=0&y=1\\ \hline\cr r_{10}&1&0\\ r_{11}&1&1\\ r_{01}&0&1\\ r_{00}&0&0\end{array}$
and therefore
$r_{\mathrm{cov}}(\theta)=p_{10,0}(\theta)+p_{11,0}(\theta)+p_{11,1}(\theta)+p_{01,1}(\theta).$
Equivalently, coverage fails on the two singleton mistakes $(r_{10},1)$ and $(r_{01},0)$ and on all abstentions $r_{00}$.
Other projections used in the paper.
The same mask formalism gives the remaining quantities used in Figure 1 and in the binary operational examples:
$\text{missed positive mass:}\qquad q_{10}(\theta)=\Pr(Y=1,\ C(X)=\{0\})=\Pr(Y=1,\ R_{\tau}(X)=r_{10})=p_{10,1}(\theta),$
$\text{hedged positive mass:}\qquad q_{11}(\theta)=\Pr(Y=1,\ C(X)=\{0,1\})=\Pr(Y=1,\ R_{\tau}(X)=r_{11})=p_{11,1}(\theta),$
$\text{abstention mass under }\pi_{\mathrm{CR}}:\qquad r_{\mathrm{abs}}(\theta;\pi_{\mathrm{CR}})=\big(p_{11,0}(\theta)+p_{11,1}(\theta)\big)+\big(p_{00,0}(\theta)+p_{00,1}(\theta)\big).$
These correspond respectively to a one-cell projection, another one-cell projection, and a two-row projection. Their binary masks are obtained immediately by placing ones on the selected cells.
G.4 Ratios of projections (conditional diagnostics)
Some diagnostics are conditional probabilities and therefore ratios of linear sums. For example, the purity of singleton-$1$ outputs under $\pi_{\mathrm{SI}}$ is
$\mathrm{Purity}_{1}(\theta):=\Pr\!\big(Y=1\mid C(X)=\{1\}\big)=\frac{\Pr(Y=1,\ C(X)=\{1\})}{\Pr(C(X)=\{1\})}=\frac{p_{01,1}(\theta)}{p_{01,0}(\theta)+p_{01,1}(\theta)}.$
Audit computation.
All entries of $P(\theta)$ are estimated from region–label counts on the audit set. Linear KPIs are computed by summing selected cells; conditional diagnostics are computed as ratios of such sums.
Appendix H Tox21 supplementary details
This appendix provides supplementary documentation for the Tox21 experiments reported in Section 5.2. The purpose of this appendix is reproducibility and contextualization rather than extension of results or additional claims.
H.1 Tox21 dataset
The Tox21 benchmark consists of binary toxicity outcomes for twelve biological assays spanning nuclear receptor (NR) signaling and cellular stress response (SR) pathways. Each compound is labeled as active or inactive per assay. Labels are sparse and highly imbalanced, particularly for nuclear receptor targets.
The full dataset composition and aggregated coverage summary used in this study is reported in the main text (Table 2).
H.2 Representative endpoint-level operational summaries
The SR-MMP endpoint summary is reported in the main text (Table 4). We keep one additional representative endpoint (NR-AR) here to document a complementary low-prevalence regime.
Table 8: NR-AR endpoint. Operational performance summary for the NR-AR endpoint. All reported quantities are joint probabilities normalized by the total test-set size. Rows report the singleton rate, doublet rate, and wrong-singleton rate ($P(Y=c,\,|S|=1,\,\hat{y}\neq Y)$) by true class. "point estimate $\mathcal{D}_{1}$" and "LOO 95% PI $\mathcal{D}_{1}$" denote leave-one-out point estimates and their exact beta–binomial predictive intervals computed on the calibration data. "Observed $\mathcal{D}_{2}$" reports empirical rates on an independent hold-out set. "$B$ 95% PI $\mathcal{D}_{2}$" gives beta–binomial predictive intervals for the corresponding $\mathcal{D}_{2}$ quantities, accounting for calibration uncertainty and finite test size.
Operational quantity Class point estimate $\mathcal{D}_{1}$ LOO 95% PI $\mathcal{D}_{1}$ Observed $\mathcal{D}_{2}$ BB 95% PI $\mathcal{D}_{2}$
Singleton rate Class 0 $0.163$ $[0.125,0.205]$ $0.133$ $[0.112,0.157]$
Class 1 $0.026$ $[0.011,0.047]$ $0.029$ $[0.019,0.041]$
Doublet rate Class 0 $0.796$ $[0.751,0.838]$ $0.818$ $[0.792,0.843]$
Class 1 $0.015$ $[0.005,0.031]$ $0.020$ $[0.012,0.030]$
Wrong-singleton rate Class 0 $0.086$ $[0.058,0.119]$ $0.060$ $[0.045,0.077]$
Class 1 $0.002$ $[0.000,0.010]$ $0.001$ $[0.000,0.004]$
Because conformal calibration is performed in a class-conditional (Mondrian) manner, the effective calibration size for the positive class governs feasibility and discretization effects. Several assays operate with fewer than 100 positive calibration samples, placing them in the small-sample regime targeted by SSBC.
H.3 Representation and model protocol
All Tox21 experiments use a fixed, task-agnostic molecular representation (descriptors plus Morgan fingerprints) and a fixed CatBoost training protocol (1000 iterations, depth 6, learning rate 0.1, log-loss), with no assay-specific feature engineering, class reweighting, or resampling. Molecules are parsed and sanitized using standard cheminformatics tooling, and compounds with invalid features are excluded. This intentionally non-optimized setup isolates conformal calibration and operational-envelope behavior from model-architecture tuning; observed AUROC values (approximately 0.80–0.85) are consistent with baseline Tox21 reports.
H.4 Conformal calibration and evaluation
All conformal predictors are calibrated in a class-conditional (Mondrian) fashion, with separate calibration sets for the positive and negative classes. For each run, one of three calibration strategies is applied: standard split conformal, DKWM-based correction, or SSBC, targeting $(\alpha,\delta)=(0.10,0.10)$.
Calibration thresholds are computed once per run and evaluated on a held-out test set. Operational rates (singleton and doublet frequencies) are estimated using leave-one-out cross-validation on the calibration set. These LOO estimates are used to construct prediction intervals for future deployment batches of fixed size, as described in the main text.
Appendix I Solubility supplementary details
This appendix records methodological details for the solubility scenario-planning experiments in Section 5.3. We first describe the dataset, partitioning, modeling, and calibration setup, and then document the operational artifacts used to interpret the resulting Pareto front.
I.1 Dataset and label construction
We use AquaSolDB as the source of aqueous solubility measurements ($\log S$), following the curation and quality-control procedures described in the dataset reference (Sorkun et al., 2019). For model training and scoring we discretize $\log S$ into three regimes with fixed thresholds: Insoluble ($\log S<-4$), Moderate ($4\leq\log S<-2$), and Soluble ($\log S\geq-2$), defining a three-class label space $\mathcal{Y}=\{0,1,2\}$ (Kalepu and Nekkanti, 2015).
I.2 Partitioning: random training and exchangeable evaluation
We use a random split to separate model fitting from the exchangeable design required for conformal calibration and finite-window operational analysis:
  • Random train/calibration/test split. Molecules are partitioned by a stratified random split, with approximately 70% assigned to training and the remaining molecules split equally between calibration and test, preserving class proportions.
  • Exchangeable calibration/test sets. The calibration and test sets are treated as exchangeable with respect to each other and to future draws under the evaluation regime in Section 5.3.
I.3 Molecular representation
Molecules are represented using a fixed descriptor set supplemented with Morgan fingerprints (Rogers and Hahn, 2010). Representation is kept task-agnostic; no assay- or regime-specific feature engineering is performed. Molecules that fail parsing or yield invalid descriptor values are excluded.
I.4 Model and training protocol
We train a CatBoost (Prokhorenkova et al., 2018) gradient-boosted decision tree classifier on the training set with fixed hyperparameters (depth 6, learning rate 0.05, multi-class loss), up to 1000 boosting iterations with early stopping (patience 50) based on a separate hyperparameter tuning-set performance metric. After training, the predictor is frozen and treated as infrastructure; the experiments study the conformal and operational layers rather than optimizing base accuracy.
I.5 Calibration and operationalization
Uncertainty quantification uses split conformal prediction with SSBC calibration (Section 5.1.1 and Appendix C), which selects a grid index to stabilize realized coverage semantics at user-specified confidence (Vovk, 2012b). This calibration layer supports the finite-window predictive-envelope constructions used for operational planning.
Although the classifier is trained on three classes (Insoluble, Moderate, Soluble), the planning analysis in Section 5.3 uses a binary operational objective obtained by merging Moderate and Soluble into a single Soluble class. Conformal prediction sets are computed in the binary label space $\{\mathrm{Insol},\mathrm{Sol}\}$ and operational quantities (loss, waste, hedging, decisiveness) are computed with respect to this binary interface.
I.6 Deployment-matched calibration via chemical tribes
Scenario planning conditions exchangeability on a deployment-defining event by restricting calibration to a chemically defined subpopulation.
Tribe construction.
We define coarse chemical “tribes” using RDKit $\mathrm{MolLogP}$:
[…]
We compute summary statistics of nonconformity and class probabilities stratified by tribe on both the calibration pool and the held-out test window, and use these diagnostics to select a focal deployment regime.
Focal deployment scenario.
Section 5.3 focuses on a lipophilic deployment regime. SSBC calibration is therefore performed on the restricted sample
$\mathcal{D}_{\mathrm{cal}}^{\mathrm{lip}}=\{(X_{i},Y_{i})\in\mathcal{D}_{\mathrm{cal}}:\mathrm{MolLogP}(X_{i})>3.5\},$
treated as exchangeable with respect to the intended deployment distribution.
Scenario-conditional exchangeability assumption.
This restriction is scenario conditioning, not a general covariate-shift correction. We assume deployment draws satisfy the same scenario event $E=\{\mathrm{MolLogP}(X)>3.5\}$ so calibration and deployment are exchangeable conditional on $E$. If deployment violates $E$, or if exchangeability fails within $E$, conformal validity is not guaranteed. Shift-robust validity methods are complementary and outside scope; see, e.g., Fannjiang et al. (2022).
Interpretation.
Conditioning trades calibration sample size for improved deployment match; the resulting loss of statistical efficiency appears directly as stronger SSBC corrections and wider predictive envelopes, consistent with transparent planning under finite-sample uncertainty.
I.7 Operational outcome definitions
Operational outcomes are defined in the main text via the binary prediction-set interface (Table 5). All reported quantities in the scenario-planning analysis are joint rates over the set–label outcomes P(Y=y,C=c)P(Y=y,\,C=c), scaled to expected counts over a window of $m=1000$ molecules.
I.8 Pareto-front artifact and representative regimes
With the methodological setup above fixed, Section 5.3 produces a deduplicated Pareto front in KPI space by sweeping SSBC calibration settings and retaining nondominated operating points. Table 9 shows two representative Pareto regimes drawn from that front to illustrate contrasting planning postures: a loss-minimizing regime (low irreversible exclusion of soluble compounds) and a decisiveness-maximizing regime (low total hedging). Counts are expected events per $1000$ molecules; brackets denote the corresponding predictive envelopes.
Table 9: Two representative Pareto-optimal operating regimes for the solubility scenario, illustrating contrasting planning postures. Rates are expected counts per 1000 molecules, with predictive envelopes in brackets.
Loss-minimizing regime High-decisiveness regime
$\alpha_{0}$ 0.125 0.150
$\delta_{0}$ 0.150 0.150
$\alpha_{1}$ 0.050 0.150
$\delta_{1}$ 0.075 0.100
Rate Interval Rate Interval
Loss rate 3 [0, 35] 12 [0, 58]
Waste rate 70 [27, 130] 13 [0, 49]
Total hedge rate 785 [663, 896] 514 [376, 675]
Correct soluble singleton rate 7 [0, 38] 88 [40, 155]
Correct insoluble singleton rate 135 [68, 215] 282 [196, 375]
Decisiveness ($1-\text{hedge}$) 215 486
I.9 Functional roles of calibration parameters
The Pareto sweep in Section 5.3 varies the four nominal parameters $(\alpha_{0},\delta_{0},\alpha_{1},\delta_{1})$ (with SSBC mapping them to effective deployed grid levels). Empirically the knobs are not interchangeable: changes reallocate mass among outcome categories (loss, waste, hedging, decisiveness) along channels constrained by the underlying threshold geometry.
Throughout, class 0 denotes insoluble and class 1 denotes soluble. The qualitative roles below summarize matched Pareto-optimal solutions where one parameter varies while the others are held fixed; rates are computed from the provided Pareto-front CSV and reported as events per 1000 molecules.
Role of $\alpha_{0}$: global conservatism against irreversible loss.
Decreasing $\alpha_{0}$ suppresses loss by expanding hedging; increasing $\alpha_{0}$ collapses ambiguity into more decisive singleton predictions. In many matched segments, the dominant effect is redistribution between hedged and singleton outcomes with loss already near saturation.
Role of $\delta_{0}$: fine-scale sharpening on the insoluble side.
For fixed $(\alpha_{0},\alpha_{1},\delta_{1})$, $\delta_{0}$ tends to act locally, shifting mass between hedged and singleton insoluble predictions with limited impact on loss.
Role of $\alpha_{1}$: boundary-controlled tolerance with nonlocal effects.
Although $\alpha_{1}$ is nominally associated with the soluble side, changing $\alpha_{1}$ can move the operating point across geometric boundaries, inducing nonlocal reallocations that may appear most strongly in insoluble outcomes.
Role of $\delta_{1}$: systematic hedge-to-insoluble reassignment.
Across stable regions of the front, increasing $\delta_{1}$ often converts a portion of hedged mass into insoluble singletons, reflecting how the normalized score constraints and threshold geometry position the hedging interval.
Geometric interpretation.
Overall, the Pareto front is shaped by geometric coupling rather than independent tuning: each knob reallocates probability mass along constrained channels determined by the threshold partition and the finite-sample SSBC adjustment.
I.10 Interpreting $\alpha$ versus $\delta$ on the Pareto front
Within SSBC the deployed threshold corresponds to an adjusted effective level $\tilde{\alpha}$ determined jointly by $(\alpha,\delta)$. Thus $\alpha$ and $\delta$ are best viewed as two parameterizations of a single underlying control (the effective threshold), with different practical resolution under finite-sample constraints. The approximate feasibility relation $\delta\approx(1-\alpha)^{N}$ implies
$\log\delta=N\log(1-\alpha),$
so changes in $\delta$ correspond to fine-grained (approximately logarithmic) adjustments in $\tilde{\alpha}$, whereas changes in $\alpha$ on a coarse grid can move the operating point across qualitatively different regions of the feasible manifold. This helps explain why $\alpha$ sweeps can trigger large reallocations among loss/waste/hedging outcomes, while $\delta$ sweeps more often act as local “sharpening” of decisions within a geometric regime.